### v1

# Unsupervised Contrastive Analysis for Salient Pattern Detection using Conditional Diffusion Models

Abstract
Contrastive Analysis (CA) regards the problem of identifying patterns in images that allow distinguishing between a background (BG) dataset (i.e. healthy subjects) and a target (TG) dataset (i.e. unhealthy subjects). Recent works on this topic rely on variational autoencoders (VAE) or contrastive learning strategies to learn the patterns that separate TG samples from BG samples in a supervised manner. However, the dependency on target (unhealthy) samples can be challenging in medical scenarios due to their limited availability. Also, the blurred reconstructions of VAEs lack utility and interpretability. In this work, we redefine the CA task by employing a self-supervised contrastive encoder to learn a latent representation encoding only common patterns from input images, using samples exclusively from the BG dataset during training, and approximating the distribution of the target patterns by leveraging data augmentation techniques. Subsequently, we exploit state-of-the-art generative methods, i.e. diffusion models, conditioned on the learned latent representation to produce a realistic (healthy) version of the input image encoding solely the common patterns. Thorough validation on a facial image dataset and experiments across three brain MRI datasets demonstrate that conditioning the generative process of state-of-the-art generative methods with the latent representation from our self-supervised contrastive encoder yields improvements in the generated image quality and in the accuracy of image classification. The code is available at https://github.com/CristianoPatricio/unsupervised-contrastive-cond-diff.
1 Introduction
[…]
Figure 1: Overview of the proposed contrastive-guided conditional diffusion model. The common features encoder $f_{\theta}(.)$ receives the input image $x_{0}$ to encode the latent representation $z_{bg}$, which captures common information. The decoder $p_{\theta}(x_{0}|x_{t},z_{bg})$ receives the noisy sample $x_{t}$ jointly with the common feature $z_{bg}$ to reconstruct the normal version of the input image. The saliency map is obtained by $\Delta_{SM}=\left|x_{0}-\hat{x_{0}}\right|$, where high values indicate the salient patterns. Intuitively, the class of the image can be determined based on the magnitude of $\bar{\Delta_{SM}}$.
Despite substantial progress in image classification models, they still face two significant challenges: i) dependence on extensive amounts of labelled data; and ii) limited interpretability. These challenges are particularly critical in sectors such as healthcare, where explaining the decision process is essential for clinicians to trust the model outcome, and unhealthy samples are scarce and hard to obtain Patrício et al. (2023). Contrastive Analysis (CA) represents a promising solution to this problem Zou et al. (2013); Abid et al. (2018); Weinberger et al. (2022) by learning the fundamental generative factors that distinguish a target (TG) dataset from a background (BG) dataset (referred to as salient patterns) and are shared across both datasets (referred to as common patterns). However, state-of-the-art CA methodologies have several problems: i) they often result in blurred images when utilizing Variational Autoencoders (VAEs) or are susceptible to mode collapse and unstable training when using Generative Adversarial Networks (GANs) Schlegl et al. (2019); Carton et al. (2024); ii) they depend on the availability of both healthy and unhealthy samples during training; iii) they struggle to preserve common patterns of the original image during the reconstruction process.
To address these problems, this work redefines the CA task by employing a self-supervised contrastive encoder (referred to as common features encoder) to learn a latent representation, using samples exclusively from the BG dataset during training, and approximating the salient variation factor by leveraging data augmentation techniques such as random cutout or Gaussian noise (e.g. presence of eyeglasses in CelebA or presence of tumor in brain MRIs). The problems of previous generative methods are addressed by leveraging Denoising Diffusion Probabilistic Models (DDPM) Ho et al. (2020) to learn the distribution of the BG dataset (control group) conditioned on the learned latent representation by the common features encoder. During inference, the diffusion model processes unseen samples from BG or TG distributions by substituting the salient patterns with common patterns learned during training. This enables the classification of the input image as background (normal) or target (anomalous) by assessing the magnitude of the reconstruction error.
Extensive validation on facial imagery data demonstrates improved similarity between original and reconstructed images. Additionally, comparative analysis with state-of-the-art methods for salient pattern detection for brain MRI further supports the efficacy of our approach. Our major contributions are as follows: i) a methodology for learning shared common information between background and target distributions, thus allowing the generation of a healthy (normal) version of the input image by encoding only its common information; ii) a common features encoder capable of learning input representations that are both class-invariant and instance-aligned. iii) a comprehensive evaluation of our approach on a facial imagery dataset and three brain MRI datasets, encompassing both healthy (background) and tumor (target) images.
2 Related Work
Early work in contrastive analysis (CA) relied on the use of Variational Autoencoders (VAEs) Baur et al. (2021); Behrendt et al. (2022); Louiset et al. (2023). A recent method proposed by Louiset et al. (2023), SepVAE, aims to distinguish common (healthy) from class-specific (unhealthy) patterns in image data. They utilized VAEs with regularization to encourage disentanglement between common and salient representations, along with a classification loss to separate target and background salient factors. However, the resulting reconstructions were often blurry, limiting their interpretability and utility. Alternative approaches Schlegl et al. (2019); Carton et al. (2024) leveraging Generative Adversarial Networks (GANs) have been explored, but they suffer from issues like mode collapse and unstable training. More recently, Diffusion Denoising Probabilistic Models Song et al. (2021); Preechakul et al. (2022) (DDPMs) have emerged as a promising alternative for high-quality image generation, addressing the drawbacks of both GANs and VAEs. Our work aligns closely with the setting of salient pattern detection. In the domain of brain MRI, Behrendt et al. (2024) introduced a patch-based diffusion model for discovering salient patterns in an unsupervised manner in brain MRIs. They divide the input image into predefined patches and apply noise to each patch individually in the forward pass. In the backward pass, the partly noised image is utilized to recover the noised patch. One drawback of their work is the extensive duration required for inference. Inspired by their approach, Iqbal et al. (2023) introduced masking-based regularization to train diffusion models, exploring various masking techniques applied at different levels of image modeling and frequency modeling. Nevertheless, their method also leads to increased inference time. Despite the potential of these methods for reconstruction fidelity, existing approaches struggle to preserve detailed characteristics of brain structure. To tackle this, we propose conditioning the denoising process of diffusion models with a latent representation encoding information about the common patterns of the input image, thereby preserving common information in the reconstructed output.
3 Background and Notations
3.1 Diffusion Models
Denoising diffusion probabilistic models  Ho et al. (2020) work by corrupting a training image $x_{0}\sim q(x_{0})$ with a predefined multi-step scheduled noise process to transform it into a sample from a Gaussian distribution. Then, a DNN is trained to revert the process, i.e., starting with a sample from a Gaussian distribution to generate a sample from the data distribution $q(x)$ through a sequence of $T$ sampling steps.
Forward Encoder
Given a training image $x_{0}$, the noising process consists of gradually noise-corrupting the image $x_{0}$ by adding Gaussian noise according to some variance schedule given by $\beta_{t}$:
$q(x_{t}|x_{t-1})=\mathcal{N}(x_{t};\sqrt{1-\beta_{t}}x_{t-1},\beta_{t}I).$ (1)
As shown in Song et al. (2021), the noisy version of an image $x_{0}$ at time $t$ is another Gaussian $q(x_{t}|x_{0})=\mathcal{N}(x_{t}|\sqrt{\alpha_{t}}x_{0},(1-\alpha_{t})I)$ where $\alpha_{t}=\prod_{s=1}^{t}(1-\beta_{s})$, which can be written in the form:
$x_{t}=\sqrt{\alpha_{t}}x_{0}+\sqrt{1-\alpha_{t}}\epsilon,\epsilon\sim\mathcal{N}(0,I).$ (2)
Reverse Decoder
Since the reverse process $q(x_{t-1}|x_{t})$ is intractable, it is used a DNN to approximate the distribution $p_{\theta}(x_{t-1}|x_{t})$, where $\theta$ represents the weights and biases of the network. The reverse process is then modeled using a Gaussian distribution of the form:
$p_{\theta}(x_{t-1}|x_{t})=\mathcal{N}(x_{t-1};\epsilon_{\theta}(x_{t},t),\beta_{t}I),$ (3)
where $\epsilon_{\theta}(x_{t},t)$ is a deep neural network governed by a set of parameters $\theta$. From $p_{\theta}(x_{1:T})$ in Equation 3, one can generate a sample $x_{t-1}$ from a sample $x_{t}$ via:
$x_{t-1}=\sqrt{\alpha_{t}-1}\left(\frac{x_{t}-\sqrt{1-\alpha_{t}}\epsilon_{\theta}(x_{t},t)}{\sqrt{\alpha_{t}}}\right)+\sqrt{1-\alpha_{t-1}}\epsilon_{\theta}(x_{t},t)+\sigma_{t}\epsilon_{t},$ (4)
where $\epsilon_{t}\sim\mathcal{N}(0,I)$ is standard Gaussian noise. For training the decoder, Ho et al. (2020) reformulated the variational lower bound objective function and considered the objective of predicting the total noise component added to the original image to create the noisy image at a given step. The loss function is then given by the squared difference between the predicted noise $\epsilon_{\theta}(x_{t},t)$ and the actual noise $\epsilon_{t}$, for a given time step $t$, using a U-Net Ronneberger et al. (2015):
$\mathcal{L}_{diff}=\left\|\epsilon_{\theta}(x_{t},t)-\epsilon_{t}\right\|_{2}^{2}.$ (5)
By setting $\sigma_{t}=0$ in Equation (4), the coefficient before the random noise $\epsilon_{t}$ becomes zero, resulting in the following deterministic process of generating a new sample $x_{t}$. This modification allows to speed up the sampling process without degrading the quality of the generated samples Bishop and Bishop (2024).
3.2 Contrastive Learning
Contrastive Learning (CL) approaches aim at pulling positive samples’ representations (e.g. of the same class) closer together while repelling representations of negative ones (e.g. different classes) apart from each other. Contrasting positive pairs against negative ones is an idea that dates back to previous research (Hadsell et al., 2006; Oord et al., 2019; Tian et al., 2020) and has seen various applications in different tasks, such as face recognition (Schroff et al., 2015). Let $x\in\mathcal{X}$ be an anchor sample, $x^{+}$ a positive sample (wrt to the anchor), and $x^{-}$ a negative sample. CL methods look for a parametric mapping function $f:\mathcal{X}\rightarrow\mathbb{S}^{d-1}$ that maps semantically similar samples close together in the representation space (i.e. a hypersphere) and dissimilar samples far away from each other. Once pre-trained, $f$ is fixed, and its representation is evaluated on a downstream task, such as classification, through linear evaluation on a test set. Depending on how positive and negative samples are defined, CL can be employed in self-supervised Chen et al. (2020) or supervised Khosla et al. (2020) settings.
4 Method
4.1 Common Features Encoder
The first block of our proposed approach is represented by the common features encoder. This encoder has the goal of learning an input representation which is invariant to the target variable (e.g. presence of eyeglasses in facial images or tumor on brain MRIs) but retains the common information of the input sample. The rationale of this approach is that an invariant representation can allow us to correctly reconstruct a realistic normal version of the input image, as it only encodes its common information. To explain the formulation of our encoder, we introduce two definitions specifying the properties that ensure that the encoder preserves the common information of the image.
Definition 4.1.
(Instance-aligned encoder) Given an anchor $x$, a positive sample $x^{+}$ of the same subject, and the set of negative samples $x^{-}_{j}$ (all other subjects), we say that an encoder $f$ is instance-aligned if:
$||f(x)-f(x^{-}_{j})||^{2}_{2}-||f(x)-f(x^{+})||^{2}_{2}\geq\epsilon\quad\forall j$ (6)
where $\epsilon\geq 0$. As the margin $\epsilon$ increases, $f$ will provide a better separation between different subjects. In practice, Eq. 6 can be expressed in terms of cosine similarity [^1] As representations are normalized, i.e. $||f(x)||_{2}=1$, then $sim(f(x),f(x^{+}))=1-d(f(x),f(x^{+}))$ where $d$ is a L2-distance function.: $sim(f(x),f(x^{-}_{j}))-sim(f(x),f(x^{+}))\leq-\epsilon\quad\forall j$ which corresponds to the $\epsilon$-InfoNCE loss Barbano et al. (2023):
$\mathcal{L}_{\epsilon-InfoNCE}=-\log\left(\frac{\exp(s^{+})}{\exp(s^{+}-\epsilon)+\sum_{j}\exp(s_{j}^{-})}\right)$ (7)
where $s^{+}$ and $s^{-}_{j}$ are shorthand notations for $sim(f(x),f(x^{+}))$ and $sim(f(x),f(x^{-}_{j})$ respectively. To obtain a sample $x^{+}$ of the same subject of $x$, if it is not available in the training data, it is possible to employ an augmentation scheme such as in SimCLR Chen et al. (2020), i.e. $x^{+}=t(x)$ where $t\sim\mathcal{T}$ is an augmentation operator sampled from a family of standard augmentation $\mathcal{T}$ (e.g. random transformations, cropping, etc.).
Definition 4.2.
(Class-invariant encoder) Denoting with $\mathcal{H}\subset\mathcal{X}$ the set of samples which share the same target attribute value (e.g. healthy), and assuming a binary case for simplicity, we say that $f$ is class-invariant, if $x\in\mathcal{H}$ and $x^{+}\in\mathcal{X}\setminus\mathcal{H}$. This means that the alignment in the latent space will be performed between samples with a different target attribute, hence achieving invariance. In a SSL setting, to avoid the dependence on anomalous samples, we leverage data augmentation and data manipulation techniques Dufumier et al. (2023) for approximating the distribution of the target attribute. For example, the appearance of tumors in brain MRIs can be approximated by employing random cutout, or Gaussian noise Behrendt et al. (2024) (details in Appendix B.2).
Definition 4.3.
(Common Features Encoder) An encoder $f$ preserves common patterns of the image if it is both instance-aligned and class-invariant.
The above definitions provide the theoretical support of the contrastive learning approach used for training the common features encoder. Considering that the learning process of the encoder is based on a contrastive learning strategy that is instance-aligned and class-invariant, our encoder is expected to produce features which preserve common information of the input image, regardless of whether the image belongs to the background or target set.
4.2 Conditional Diffusion-based Decoder
Our conditional diffusion decoder $p_{\theta}(x_{t-1}|x_{t},z_{bg})$ takes as input the noisy sample $x_{t}$ and the common feature $z_{bg}\in\mathbb{R}^{1\times d}=f_{\theta}(x_{0})$, a non-spatial vector of dimension $d$ that encodes common patterns observed in the input sample, derived from the properties of the common features encoder $f_{\theta}$ (Section 4.1). Our primary objective is to reconstruct the background version of the input image, preserving its common information. Hence, we deviate from the step-wise sampling process in Equation (3), which is typically employed to generate new images from noise $p_{\theta}(x_{T:1}),x_{T}\sim\mathcal{N}(0,I)$. Instead, we directly estimate the input image $\hat{x_{0}}\sim p_{\theta}(x_{0}|x_{t},z_{bg})$ given the noisy sample $x_{t}$ (more details in Algorithm 2 in Appendix A.2). This is achieved by revising Equation (2), following Song et al. (2021), enabling to prediction of the denoised observation, which is an estimation of $x_{0}$ given $x_{t}$:
$g_{\theta}(x_{t},t,z_{bg}):=(x_{t}-\sqrt{1-\alpha_{t}}\cdot\epsilon_{\theta}(x_{t},t,z_{bg}))/\sqrt{\alpha_{t}}$ (8)
Then, the model is trained (Algorithm 1 in Appendix A.1) by minimizing the following objective function Preechakul et al. (2022), which is a modified version of the MSE objective in Equation (5):
$\mathcal{L}_{diff}=\sum_{t=1}^{T}\mathbb{E}_{x_{0},\epsilon_{t}}\left[\left\|\epsilon_{\theta}(x_{t},t,z_{bg})-\epsilon_{t}\right\|_{2}^{2}\right]$ (9)
We utilize adaptive group normalization layers (AdaGN) to condition the U-Net, as introduced in prior works Dhariwal and Nichol (2021); Preechakul et al. (2022). AdaGN incorporates the timestep and conditional variable embedding into each residual block by applying channel-wise scaling and shifting on the normalized intermediate activations $h\in\mathbb{R}^{c\times h\times w}$: $\textrm{AdaGN}(h,t,z_{bg})=z_{i}(t_{s}\textrm{GroupNorm}(h)+t_{b})$, where $z_{i}\in\mathbb{R}^{c}=\textrm{Affine}(z_{bg})$ and $(t_{s},t_{b})\in\mathbb{R}^{2\times c}=MLP(\psi(t))$ is the output of a multilayer perceptron incorporating a sinusoidal encoding function $\psi$. Additional details are provided in Appendix A.
5 Experiments
5.1 Experimental Data
Facial Images
We conduct initial experiments on CelebA Liu et al. (2015), containing 202,599 facial images with diverse attributes. We create a subset focusing on subjects wearing eyeglasses and those without accessories, resulting in 15,353 images divided into two distinct classes: 1) Eyeglasses (EG): Images with the ’Eyeglasses’ attribute and no other accessory-related attributes, used solely for evaluation as target images; 2) No Eyeglasses (NEG): Images without the ’Eyeglasses’ attribute or any accessory-related attributes, used for training as background images. More information on dataset partitions and pre-processing is provided in Appendices B.1 and B.2.
Medical Images
To evaluate our common features encoder on medical data, we focus on tumor detection in brain MRIs. For a comprehensive comparison, we utilize the IXI IXI dataset as a background reference for training, as done in previous studies Behrendt et al. (2024); Iqbal et al. (2023). Evaluation is conducted on the Multimodal Brain Tumor Segmentation Challenge 2021 (BraTS21) Menze et al. (2014); Bakas et al. (2017); Baid et al. (2021) and the multiple sclerosis dataset (MSLUB) Lesjak et al. (2018), featuring tumor and Multiple Sclerosis (MS) samples, respectively. Notably, only T2-weighted images are used from all datasets. More information on preprocessing and data partitioning is provided in Appendix B.3.
5.2 Experimental Setup
Implementation Details
We compare our approach with state-of-the-art diffusion-based methods on CelebA Liu et al. (2015), including Diff-AE Preechakul et al. (2022), SepVAE Louiset et al. (2023), and DDIM Dhariwal and Nichol (2021). To ensure fair comparison, we use the official implementation of these methods and train them on our dataset partitions. For MRI experiments, we solely utilize our common features encoder alongside DDPM Wyatt et al. (2022), $p$DDPM Behrendt et al. (2024), and $m$DDPM Iqbal et al. (2023). These methods adopt distinct noise and objective functions, indirectly comparable to DDIM in facial images. Notably, they use structured simplex noise instead of Gaussian noise, capturing MRI image frequency distribution better. Additionally, they employ an alternative objective function to directly minimize $l1$ error between input image $x_{0}$ and its reconstruction $\hat{x_{0}}$ (Equation (9)). During training, MRI volumes are processed slice-wise, uniformly sampled with replacement. At test time, full volume reconstruction is achieved by iterating over all slices [^2] Implementation code of our method will be publicly available..
Evaluation Pipeline
During inference, the salient pattern map is obtained by the absolute difference between the input image $x_{0}$ and the reconstructed image $\hat{x_{0}}$, where higher values indicate larger reconstruction errors. To enhance the quality of these maps, inspired by Behrendt et al. (2024), we apply a median filter with a kernel size of 5 for score smoothing, followed by a morphological erosion operation for 10 iterations. Subsequently, the map is binarized, and the threshold is fine-tuned on the validation set to maximize the average DICE score. Additionally, we calculate the average Area Under the Precision-Recall Curve (AUPRC) on the test set.
5.3 Results
Table 1: Reconstruction quality of methods trained on CelebA (training set) and tested on CelebA (test set). NEG - No Eyeglasses, EG - Eyeglasses. $\left|\Delta\right|=\left|NEG-EG\right|$.
Model SSIM $\uparrow$ MSE $\downarrow$ LPIPS $\downarrow$
NEG EG $\left|\Delta\right|$ NEG EG $\left|\Delta\right|$ NEG EG $\left|\Delta\right|$
Diff-AE (No $\textrm{x}_{T}$)  Preechakul et al. (2022) $0.6287$ $0.6079$ $0.0208$ 0.01040.0104{0.0104}0.0104 $0.0119$ $0.0015$ $0.0847$ 0.10410.1041{0.1041}0.1041 $0.0194$
SepVAE Louiset et al. (2023) $0.5796$ $0.4773$ $0.1023$ $0.0153$ $0.0268$ $0.0115$ $0.2912$ $0.3322$ $0.0410$
DDIM Dhariwal and Nichol (2021) 0.96900.9690{0.9690}0.9690 $0.6568$ 0.3122 $\underline{0.0004}$ $0.0095$ $0.0091$ 0.00770.0077{0.0077}0.0077 $0.1385$ 0.1308
Ours (w/o RE) $\underline{0.9756}$ $0.6737$ $0.3019$ 0.0003 $0.0097$ $0.0094$ $\underline{0.0060}$ $0.1198$ $0.1138$
Ours (w/ RE) 0.9763 $0.6701$ $\underline{0.3062}$ 0.0003 $0.0099$ 0.0096 0.0059 $0.1213$ $\underline{0.1154}$
5.3.1 CelebA Dataset
Table 1 presents the reconstruction quality results of our method alongside Diff-AE Preechakul et al. (2022), SepVAE Louiset et al. (2023), and DDIM Dhariwal and Nichol (2021). These models were trained on the CelebA dataset using only background samples (NEG) and evaluated on both background (NEG) and target (EG) images. Evaluation metrics include SSIM Wang et al. (2004) ($\uparrow$), LPIPS Zhang et al. (2018) ($\downarrow$), and MSE ($\downarrow$). Our method (w/ RE) achieves the highest SSIM (0.9763) and lowest MSE (0.0003) and LPIPS (0.0059) when reconstructing background examples (NEG). While our method (w/o RE) and DDIM produce similar results, they fail to capture common information and high-level details compared to our method (w/ RE), as visually demonstrated in Figure 2. Additionally, ablation studies investigate the reconstruction quality when different $t\in[1,T]$ with $T=1000$ are chosen for corrupting the input image to predict its denoised version (Figure 10 in Appendix D). Empirically, we find that $t=250$ strikes a balance between image quality and common information preservation.
EG-NEG Classification Performance
Table 1 provides metrics for distinguishing between the two classes (NEG vs. EG), with the magnitude $\left|\Delta\right|=\left|NEG-EG\right|$ indicating distinctiveness. Histogram plots are available in Figure 7 in Appendix C. Diff-AE Preechakul et al. (2022) and SepVAE Louiset et al. (2023) exhibit overlapping distributions across metrics, while our method demonstrates well-separated distributions, yielding the highest ROC AUC score (Figure 8 in Appendix C).
Instance-alignment
Varying the margin $\epsilon$ in Equation (7) can influence identification accuracy by affecting the separation of different classes within the latent space. We conduct an ablation study on $\epsilon$ to observe its impact on identification accuracy using a $k$-NN classifier with $k=3$. Results depicted in Figure 3 (left) show that the highest accuracy is achieved when $\epsilon$ is chosen from the ranges $[0.5,2.5]$ with the RE transformation and $[0.15,1.0]$ without it.
Importance of class-invariance
We conduct a linear probing analysis to assess the encoder’s class-invariance (Definition 4.2). The objective is to evaluate the encoder’s capability to distinguish between EG and NEG samples. Subsequently, a logistic regression classifier was trained on the extracted image latent features. Results in Figure 3 show lower target accuracy with RE, suggesting stronger common information preservation. However, instance classification performs worse with RE, possibly due to built-in augmentations in contrastive learning. Implementation details and numerical results are available in Appendix D.2.
OriginalDiff-AESepVAEDDIMOurs(w/o RE)Ours(w/ RE)OriginalDiff-AESepVAEDDIMOurs(w/o RE)Ours(w/ RE)Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
Figure 2: Reconstruction results of CelebA images using different methods. In contrast to Diff-AE, SepVAE and DDIM, our method produces images with well-preserved common information from the input image.
$\epsilon$$\epsilon$Refer to captionRefer to caption
Figure 3: Left: Instance classification (accuracy in %) for different values of $\epsilon$. Right: Classification accuracy (NEG vs. EG) for different values of $\epsilon$.
Table 2: Comparison of the evaluated models on brain MRI datasets. The best results are highlighted in bold, and the second-best results are underlined. * denotes that the results were obtained by reproducing the method. The coloured values between brackets indicate the improvement w.r.t. the correspondent method. The reported results for all metrics correspond to the mean $\pm$ standard deviation across the different folds.
Model BraTS21 MSLUB IXI
DICE (%)\;$\uparrow$ AUPRC (%)\;$\uparrow$ DICE (%)\;$\uparrow$ AUPRC (%)\;$\uparrow$ $l1(1e-3)\downarrow$
AE  Baur et al. (2021) $32.87$±$\pm 1.25$ $31.07$±$\pm 1.75$ $7.10$±$\pm 0.68$ $5.58$±$\pm 0.26$ $30.55$±$\pm 0.27$
VAE Baur et al. (2021) $31.11$±$\pm 1.50$ $28.80$±$\pm 1.92$ $6.89$±$\pm 0.09$ $5.00$±$\pm 0.40$ $31.28$±$\pm 0.71$
SVAE Behrendt et al. (2022) $33.32$±$\pm 0.14$ $33.14$±$\pm 0.20$ $5.76$±$\pm 0.44$ $5.04$±$\pm 0.13$ $28.08$±$\pm 0.02$
DAE Kascenas et al. (2022) $37.05$±$\pm 1.42$ 44.9944.99{44.99}44.99±$\pm 1.72$ $3.56$±$\pm 0.91$ $5.35$±$\pm 0.45$ $10.12$±$\pm 0.26$
f-AnoGAN Schlegl et al. (2019) $24.16$±$\pm 2.94$ $22.05$±$\pm 3.05$ $4.18$±$\pm 1.18$ $4.01$±$\pm 0.90$ $45.30$±$\pm 2.98$
DDPM* Wyatt et al. (2022) $39.25$±$\pm 1.01$ $47.79$±$\pm 1.28$ $5.43$±$\pm 1.71$ $7.25$±$\pm 0.79$ $14.10$±$\pm 1.64$
pDDPM* Behrendt et al. (2024) $49.47$±$\pm 0.91$ $54.68$±$\pm 1.02$ $9.17$±$\pm 1.29$ $\underline{10.35}$±$\pm 0.80$ $11.31$±$\pm 0.91$
mDDPM* Iqbal et al. (2023) $\underline{51.77}$±$\pm 0.18$ $\underline{57.49}$±$\pm 0.41$ $5.03$±$\pm 0.37$ $9.47$±$\pm 0.02$ 7.287.28{7.28}7.28±$\pm 0.00$
Ours (Common Features Encoder $+$ DDPM) 39.25±$\pm 0.55$ (+0.00) 48.58±$\pm 1.06$ (+0.79) $5.39$±$\pm 1.13$ 7.29±$\pm 0.23$ (+0.04) 13.72±$\pm 0.56$ (+0.38)
Ours (Common Features Encoder $+$ pDDPM) $49.36$±$\pm 0.53$ $54.60$±$\pm 0.60$ 10.48±$\pm 2.19$ (+1.31) 10.2610.26{10.26}10.26±$\pm 0.67$ 11.11±$\pm 0.01$ (+0.20)
Ours (Common Features Encoder $+$ mDDPM) $51.69$±$\pm 0.17$ 57.58±$\pm 0.17$ (+0.09) 9.47±$\pm 0.18$ (+4.44) $9.44$±$\pm 0.48$ $7.47$±$\pm 0.00$
Identity-conditional sampling
To demonstrate the efficacy of our encoder in capturing background patterns from a query image $x_{0}\sim q(x_{0})$, we encode $x_{0}$ to derive its latent representation $z_{bg}=f_{\theta}(x_{0})$. Then, using the step-wise denoising process $p_{\theta}(x_{t-1}|x_{t},z_{bg})$, we generate new images sharing the common information of $x_{0}$. Examples are shown in Figure 9 in Appendix C. With our encoder (w/ RE), the generated images preserve most background features and facial pose. Notably, subjects’ eyes appear more natural and perceptible with RE, suggesting the encoder is class-invariant, focusing on common patterns between the query and generated images.
5.3.2 Brain MRI Datasets
Table 2 provides a comparison between different unsupervised salient pattern detection methods in brain MRI. Evaluation metrics include DICE score, AUPRC, and $l1$ reconstruction error on healthy data (IXI). We reproduced results for DDPM Wyatt et al. (2022), pDDPM Behrendt et al. (2024), and mDDPM Iqbal et al. (2023). Results for other methods are referenced from the work of Behrendt et al. Behrendt et al. (2024).
Reconstruction Results
As depicted in Table 2, incorporating our common features encoder into diffusion-based models results in slight improvements across all datasets. Particularly noteworthy is the enhancement observed over the base results for the IXI dataset when our encoder is utilized. This suggests that when reconstructing a background (healthy) sample, the generated image preserves most of the common patterns from the input image. Detailed reconstruction examples can be found in Figure 11 in Appendix D.
Salient Pattern Detection
Examples of saliency maps generated by various methods are illustrated in Figure 4. Notably, we observe reduced foreground noise in the maps generated by $m$DDPM compared to DDPM, particularly evident in regions representing healthy brain anatomy. Specifically, when comparing the maps of the image depicted in the bottom-right row, the combination of Ours+DDPM and Ours+$m$DDPM demonstrates a smaller reconstruction error.
[…]
Figure 4: Salient pattern maps generated by different methods. Lighter colors indicate larger anomalies.
6 Conclusion and Future Work
In this work, we introduce a self-supervised contrastive encoder, termed the common features encoder, designed to learn a latent representation capturing common patterns from input images. We then use diffusion-based models conditioned on this representation to generate realistic images exclusively encoding these common patterns. The encoder is trained to ensure class invariance and instance alignment, enabling the effective capture of common information during the reconstruction process. Our encoder demonstrates notable effectiveness in the CelebA dataset, and when integrated into established diffusion models for brain MRI, it improves the baseline results by preserving more background (common) patterns in the reconstructed images. Furthermore, investigating alternative mechanisms to integrate the latent condition into diffusion model architecture is a promising future direction. The potential implications of our approach on clinical workflows are discussed in Appendix E.
Appendix A Conditional Diffusion Model Architecture
The contrastive-guided conditional diffusion model is based on the DDIM Dhariwal and Nichol [2021] model, whereas the incorporation of the condition $z_{bg}$ into the U-Net model is based on the Diff-AE Preechakul et al. [2022] model. We selected the hyperparameters according to our computational resources. Figure 5 illustrates the operations inside the ResBlock to incorporate both the timestep $t$ and the condition $z_{bg}$, where $MLP$ denotes an embedding layer applied to both the timestep embedding and the condition $z_{bg}$ to match the number of output channels of the previous layer.
Refer to caption $MLP(\psi(t))$ $MLP(z_{bg})$ $MLP(\psi(t))$ $MLP(z_{bg})$
Figure 5: Diagram architecture of our approach. The bottom image illustrates a Residual Block along with the internal operations to include the timestep $t$ and the condition $z_{bg}$.
A.1 Model Training
The algorithm for model training is detailed in Algorithm 1.
Algorithm 1 Conditional Diffusion Learning
Input: Training data $\mathcal{D}=\{x_{n}\}$, Noise schedule $\{\beta_{1},...,\beta_{T}\}$
Output: Network parameters w
for $t\in\{1,...,T\}$ do
   $\alpha_{t}\leftarrow\prod_{\tau=1}^{t}(1-\beta_{\tau})$
end for
repeat
   $x\sim\mathcal{D}$
   $t\sim\{1,...,T\}$
   $z_{bg}\leftarrow f_{\theta}(x)$
   $\epsilon\sim\mathcal{N}(0,I)$
   ${x}_{t}\leftarrow\sqrt{\alpha_{t}}{x}+\sqrt{1-\alpha_{t}}\epsilon$
   $\mathcal{L}(\textbf{w})\leftarrow\mathcal{L}_{diff}$
   Take optimizer step
until converge
return w
A.2 Model Inference
The algorithm for predicting the normal (healthy) version of the input image is detailed in Algorithm 2.
Algorithm 2 Denoising Process for Normal Image Inference
Input: Test data $\mathcal{D}=\{x_{n}\}$, Noise schedule $\{\beta_{1},...,\beta_{T}\}$, Trained denoising network $\epsilon_{\theta}(x_{t},t,z_{bg})$
Output: Denoised observation $\hat{x_{0}}$
for $t\in\{1,...,T\}$ do
   $\alpha_{t}\leftarrow\prod_{\tau=1}^{t}(1-\beta_{\tau})$
end for
$t\sim\{1,...,T\}$
$\epsilon\sim\mathcal{N}(0,I)$
${x}_{t}\leftarrow\sqrt{\alpha_{t}}{x}+\sqrt{1-\alpha_{t}}\epsilon$
$z_{bg}\leftarrow f_{\theta}(x)$
$\hat{x_{0}}\leftarrow g_{\theta}(x_{t},t,z_{bg})$
return $\hat{x_{0}}$
Appendix B Data Preprocessing
B.1 CelebA Dataset
The images per each split were selected based on the available official partitions and filtered based on the attribute annotations Liu et al. [2015]. We filtered the CelebA dataset to obtain images containing the attribute Eyeglasses and images without any attribute related to wearing accessories. Specifically, we obtain the images of subjects with Eyeglasses by searching for the images matching the following attributes: Eyeglasses $=1$, Wearing_Hat $=-1$, Wearing_Earrings $=-1$, Wearing_Lipstick $=-1$, Wearing_Necklace $=-1$, Wearing_Necktie $=-1$. Detailed statistics of the filtered CelebA, including the train/validation/test partitions, are presented in Table 3.
Table 3: CelebA Dataset Statistics.
Split Train Validation Test
Class NEG EG NEG EG NEG EG -
# 6488 5931 800 764 712 658 -
Total 12419 1564 1370 15353
B.2 Custom Random Erasing
We modify the RandomErasing class Zhong et al. [2020] available in the torchvision.transforms to include a parameter $b$ that controls the level of random noise (governed by a Gaussian distribution $\mathcal{N}(0,1-b),b\in[0,1]$) added to the rectangle region with respect to its spatial location and aspect ratio, where $b=1$ means that no noise is added. Figure 6 illustrates the augmented versions of the images (from CelebA and IXI datasets) after applying our custom random-erasing during the encoder training.
Refer to caption Refer to caption
(a) CelebA (b) IXI
Figure 6: (a) - Left column: anchor image. Right column: augmented image after applying our custom random-erasing transformation. (b) - Left column: anchor image. Right column: augmented image after applying our custom random-erasing transformation.
B.3 MRI Datasets
  • IXI IXI dataset comprises 560 pairs of T1 and T2-weighted brain MRI scans. From the 560 pairs, 158 samples are used for testing and the remaining data is divided into 5 folds of 358 training samples and 44 validation samples for cross-validation.
  • BraTS21 Menze et al. [2014], Bakas et al. [2017], Baid et al. [2021] dataset comprises 1251 brain MRI scans of four different weightings (T1, T1-CE, T2, FLAIR). The data is split into an unhealthy validation set of 100 samples and an unhealthy test set of 1151 samples.
  • MSLUB Lesjak et al. [2018] dataset comprises brain MRI scans of 30 patients with Multiple Sclerosis (MS). For each patient, T1, T2, and FLAIR-weighted scans are available. The data is split into an unhealthy validation set of 10 samples and an unhealthy test set of 20 samples.
Following Behrendt et al. [2024], we utilized T2-weighted images from all datasets across our experiments, as well as performed some pre-processing steps involving brain registration, skull stripping and downsampling. Please refer to Behrendt et al. [2024] for more detailed information on this pre-processing steps.
Appendix C Additional Results
Figure 7 shows the histograms plots of class-label distributions amongst MSE, SSIM and LPIPS on CelebA for the evaluated methods (Diff-AE, SepVAE and Ours).
(a) Diff-AE Preechakul et al. [2022]
(b) SepVAE Louiset et al. [2023]
(c) Ours (w/ RE)
Figure 7: Histogram of MSE, SSIM, and LPIPS for CelebA for NEG and EG samples.
The ROC curves on CelebA for MSE, SSIM and LPIPS are illustrated in Figure 8.
Refer to caption Refer to caption Refer to caption
(a) MSE (b) SSIM (c) LPIPS
Figure 8: ROC curve on CelebA for (a) MSE, (b) SSIM and (c) LPIPS.
Figure 9 depicts examples of images generated by DDIM using the setp-wise denoining process $p_{\theta}(x_{t-1}|x_{t},z_{bg})$, conditioned on the common latent feature $z_{bg}$.
QueryOurs (w/ RE)Ours (w/o RE)Close-upRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
Figure 9: Identity-conditional generation using the latent representation $z_{bg}$ of the query image. Incorporating random-erasing (RE) during encoder training encourages class-invariance (Section 4.1), as demonstrated by the close-up on the eyes region, where w/o RE the eyes appear uniformly occluded.
Appendix D Ablation Studies
D.1 Sampling Facial Images at Different Timesteps
Figure 10 shows the results for three facial images. We noticed that when choosing a $t>350$, the image gradually loses background (common) information. Conversely, $t<100$ fails to effectively replace eyeglasses, especially if they are sunglasses occluding the eyes (subject of first row). However, for normal eyeglasses without occluding the eyes (subject of last row), $t=100$ proves to be an optimal choice.
[…]
Figure 10: Predicted $\textrm{x}_{0}$ at different time steps. The predicted image gradually looses details and is altered as the level of noise increases.
D.2 Instance-alignment and Class-invariant of Common Features Encoder
To assess the influence of $\epsilon$ on both subject identification accuracy and EG/NEG classification accuracy, we conducted an ablation study on $\epsilon$ to verify the instance-alignment and class-invariance properties of our encoder (Section 4.1). The results are reported in Table 4. Specifically, for instance classification, we employed a $k$-NN classifier ($k=3$) and calculated the cosine similarity between the latent representation of each test sample and the remaining test samples. For EG/NEG classification, a logistic regression classifier was trained on the extracted latent features of images to distinguish between EG and NEG examples.
Table 4: Results (Accuracy in %) of Instance Classification and EG/NEG Classification tasks on CelebA.
$\epsilon$ w/o RE w/ RE
EG/NEG Class. (%) Instance Class. (%) ($3-NN$) EG/NEG Class. (%) Instance Class. (%) ($3-NN$)
0.0 80.219 24.14 76.715 22.55
0.25 84.599 24.73 74.307 22.80
0.5 82.993 26.32 74.088 22.72
1.0 83.431 25.48 72.336 23.72
1.5 82.555 22.97 72.701 23.05
2.0 84.453 23.47 72.190 23.72
2.5 80.876 24.06 70.949 22.72
3.0 80.657 23.72 71.679 20.54
3.5 80.803 24.48 68.686 21.29
4.0 78.978 22.13 69.270 21.71
4.5 78.467 23.22 68.832 19.87
5.0 77.883 21.71 67.299 20.28
6.0 78.686 23.81 64.161 19.95
7.0 75.036 22.05 62.263 19.20
8.0 73.358 20.54 62.774 20.12
Diff-AE Preechakul et al. [2022] Target Accuracy: 87.664
D.3 Reconstructed Brain MRI images
In Figure 11, the reconstruction version of brain MRIs are depicted for the evaluated methods. Notably, the combination Ours+$p$DDPM achieved the lowest reconstruction $l1$ error. Ablation results on $\epsilon$ are present in Table 5.
[…]
Figure 11: Reconstruction results of MRI images using different methods.
Table 5: Ablation results on $\epsilon$ for the combination of our encoder with the evaluated models on brain MRI datasets. The best results are highlighted in bold, and the second-best results are underlined. * denotes that the results were obtained by reproducing the method. The coloured values between brackets indicate the improvement/decrease w.r.t. the correspondent method. The reported results for all metrics correspond to the mean $\pm$ standard deviation across the different folds.
Model BraTS21 MSLUB IXI
DICE (%)\;$\uparrow$ AUPRC (%)\;$\uparrow$ DICE (%)\;$\uparrow$ AUPRC (%)\;$\uparrow$ $l1(1e-3)\downarrow$
AE  Baur et al. [2021] $32.87$±$\pm 1.25$ $31.07$±$\pm 1.75$ $7.10$±$\pm 0.68$ $5.58$±$\pm 0.26$ $30.55$±$\pm 0.27$
VAE Baur et al. [2021] $31.11$±$\pm 1.50$ $28.80$±$\pm 1.92$ $6.89$±$\pm 0.09$ $5.00$±$\pm 0.40$ $31.28$±$\pm 0.71$
SVAE Behrendt et al. [2022] $33.32$±$\pm 0.14$ $33.14$±$\pm 0.20$ $5.76$±$\pm 0.44$ $5.04$±$\pm 0.13$ $28.08$±$\pm 0.02$
DAE Kascenas et al. [2022] $37.05$±$\pm 1.42$ 44.9944.99{44.99}44.99±$\pm 1.72$ $3.56$±$\pm 0.91$ $5.35$±$\pm 0.45$ $10.12$±$\pm 0.26$
f-AnoGAN Schlegl et al. [2019] $24.16$±$\pm 2.94$ $22.05$±$\pm 3.05$ $4.18$±$\pm 1.18$ $4.01$±$\pm 0.90$ $45.30$±$\pm 2.98$
DDPM* Wyatt et al. [2022] $39.25$±$\pm 1.01$ $47.79$±$\pm 1.28$ $5.43$±$\pm 1.71$ $7.25$±$\pm 0.79$ $14.10$±$\pm 1.64$
pDDPM* Behrendt et al. [2024] $49.47$±$\pm 0.91$ $54.68$±$\pm 1.02$ $9.17$±$\pm 1.29$ 10.35±$\pm 0.80$ $11.31$±$\pm 0.91$
mDDPM* Iqbal et al. [2023] 51.77±$\pm 0.18$ $57.49$±$\pm 0.41$ $5.03$±$\pm 0.37$ $9.47$±$\pm 0.02$ $\underline{7.28}$±$\pm 0.00$
Ours (Common Features Encoder $+$ DDPM)
$\epsilon=0.5$ $38.68$±$\pm 0.85$ (-0.57) $48.49$±$\pm 1.47$ (+0.70) $6.47$±$\pm 1.51$ (+1.04) $7.69$±$\pm 1.30$ (+0.44) $13.78$±$\pm 0.52$ (+0.32)
$\epsilon=1.0$ $39.01$±$\pm 0.60$ (-0.24) $48.01$±$\pm 1.30$ (+0.22) $5.44$±$\pm 2.06$ (+0.01) $7.97$±$\pm 0.99$ (+0.72) $13.85$±$\pm 0.60$ (+0.25)
$\epsilon=2.0$ $39.25$±$\pm 0.55$ (+0.00) $48.58$±$\pm 1.06$ (+0.79) $5.39$±$\pm 1.13$ (-0.04) $7.29$±$\pm 0.23$ (+0.04) $13.72$±$\pm 0.56$ (+0.38)
$\epsilon=3.0$ $39.21$±$\pm 0.36$ (-0.04) $48.08$±$\pm 1.57$ (+0.29) $5.35$±$\pm 1.25$ (-0.08) $7.38$±$\pm 0.81$ (+0.13) $13.79$±$\pm 0.42$ (+0.31)
Ours (Common Features Encoder $+$ pDDPM)
$\epsilon=0.5$ $48.93$±$\pm 1.42$ (-0.54) $54.06$±$\pm 1.54$ (-0.62) 10.55±$\pm 1.42$ (+1.38) $9.94$±$\pm 0.92$ (-0.41) $11.15$±$\pm 0.01$ (+0.16)
$\epsilon=1.0$ $49.26$±$\pm 0.74$ (-0.21) $54.65$±$\pm 0.86$ (-0.03) $9.29$±$\pm 2.25$ (+0.12) $9.90$±$\pm 1.22$ (-0.45) $11.13$±$\pm 0.02$ (+0.18)
$\epsilon=2.0$ $48.84$±$\pm 0.61$ (-0.63) $53.96$±$\pm 0.75$ (-0.72) $9.62$±$\pm 0.97$ (+0.45) $9.71$±$\pm 0.44$ (-0.64) $11.12$±$\pm 0.02$ (+0.19)
$\epsilon=3.0$ $49.36$±$\pm 0.53$ (-0.11) $54.60$±$\pm 0.60$ (-0.08) $\underline{10.48}$±$\pm 2.19$ (+1.31) $\underline{10.26}$±$\pm 0.67$ (-0.09) $11.11$±$\pm 0.01$ (+0.20)
Ours (Common Features Encoder $+$ mDDPM)
$\epsilon=0.5$ $\underline{51.76}$±$\pm 0.85$ (-0.01) $\underline{57.56}$±$\pm 0.96$ (+0.07) $8.66$±$\pm 1.40$ (+3.63) $9.34$±$\pm 1.24$ (-0.13) $7.36$±$\pm 0.00$ (-0.08)
$\epsilon=1.0$ $51.66$±$\pm 0.04$ (-0.11) $57.27$±$\pm 0.03$ (-0.22) $8.69$±$\pm 0.33$ (+3.66) $9.51$±$\pm 0.15$ (+0.04) 7.24±$\pm 0.79$ (+0.04)
$\epsilon=2.0$ $51.50$±$\pm 0.58$ (-0.27) $57.15$±$\pm 0.57$ (-0.34) $8.75$±$\pm 0.96$ (+3.72) $9.22$±$\pm 0.42$ (-0.25) $7.39$±$\pm 0.00$ (-0.11)
$\epsilon=3.0$ $51.69$±$\pm 0.17$ (-0.08) 57.58±$\pm 0.17$ (+0.09) $9.47$±$\pm 0.18$ (+4.44) $9.44$±$\pm 0.48$ (-0.03) $7.47$±$\pm 0.00$ (-0.19)
Appendix E Potential Impact on Clinical Workflow
The ability to reconstruct normal representations of input images, regardless of the presence of abnormalities, holds considerable significance in the disease diagnosis process. Early detection of diseases is aided by comparing the original image with its reconstructed healthy counterpart. Discrepancies between the images may signify the existence of anomalies or abnormalities, such as tumors. The development of an AI tool designed to assist healthcare practitioners in promptly identifying diseases or abnormalities by comparing two images enables timely intervention and treatment. Additionally, providing clinicians with reconstructed healthy images for comparison purposes can contribute to the reduction of diagnostic errors by providing clearer visualization of abnormalities and assisting in the process of differential diagnosis.

### v3

@@
-Unsupervised Contrastive Analysis for Salient Pattern Detection using Conditional Diffusion Models
+Unsupervised contrastive analysis for anomaly detection in brain MRIs via conditional diffusion models
 Abstract
-Contrastive Analysis (CA) regards the problem of identifying patterns in images that allow distinguishing between a background (BG) dataset (i.e.
-healthy subjects) and a target (TG) dataset (i.e.
-unhealthy subjects).
-Recent works on this topic rely on variational autoencoders (VAE) or contrastive learning strategies to learn the patterns that separate TG samples from BG samples in a supervised manner.
-However, the dependency on target (unhealthy) samples can be challenging in medical scenarios due to their limited availability.
-Also, the blurred reconstructions of VAEs lack utility and interpretability.
-In this work, we redefine the CA task by employing a self-supervised contrastive encoder to learn a latent representation encoding only common patterns from input images, using samples exclusively from the BG dataset during training, and approximating the distribution of the target patterns by leveraging data augmentation techniques.
-Subsequently, we exploit state-of-the-art generative methods, i.e.
-diffusion models, conditioned on the learned latent representation to produce a realistic (healthy) version of the input image encoding solely the common patterns.
-Thorough validation on a facial image dataset and experiments across three brain MRI datasets demonstrate that conditioning the generative process of state-of-the-art generative methods with the latent representation from our self-supervised contrastive encoder yields improvements in the generated image quality and in the accuracy of image classification.
-The code is available at https://github.com/CristianoPatricio/unsupervised-contrastive-cond-diff.
+Contrastive Analysis (CA) detects anomalies by contrasting patterns unique to a target group (e.g., unhealthy subjects) from those in a background group (e.g., healthy subjects).
+In the context of brain MRIs, existing CA approaches rely on supervised contrastive learning or variational autoencoders (VAEs) using both healthy and unhealthy data, but such reliance on target samples is challenging in clinical settings.
+Unsupervised Anomaly Detection (UAD) offers an alternative by learning a reference representation of healthy anatomy without the need for target samples.
+Deviations from this reference distribution can indicate potential anomalies.
+In this context, diffusion models have been increasingly adopted in UAD due to their superior performance in image generation compared to VAEs.
+Nonetheless, precisely reconstructing the anatomy of the brain remains a challenge.
+In this work, we propose an unsupervised framework to improve the reconstruction quality by training a self-supervised contrastive encoder on healthy images to extract meaningful anatomical features.
+These features are used to condition a diffusion model to reconstruct the healthy appearance of a given image, enabling interpretable anomaly localization via pixel-wise comparison.
+We validate our approach through a proof-of-concept on a facial image dataset and further demonstrate its effectiveness on four brain MRI datasets, achieving state-of-the-art anomaly localization performance on the NOVA benchmark.
+keywords:
+Contrastive Analysis , Unsupervised Anomaly Detection , Brain MRI , Contrastive Learning , Neuroimaging , Diffusion Models
+journal: Pattern Recognition Letters
+[…]
+[ubi] organization=Universidade da Beira Interior and NOVA LINCS,city=Covilhã, country=Portugal
+[…]
+[feup] organization=Faculdade de Engenharia da Universidade do Porto,city=Porto, country=Portugal
+[…]
+[unito] organization=University of Turin,city=Turin, country=Italy
+[…]
+[inesctec] organization=INESC TEC,city=Porto, country=Portugal
 1 Introduction
-[…]
 Figure 1: Overview of the proposed contrastive-guided conditional diffusion model.
-The common features encoder $f_{\theta}(.)$ receives the input image $x_{0}$ to encode the latent representation $z_{bg}$, which captures common information.
+The target-invariant encoder $f_{\theta}(.)$ receives the input image $x_{0}$ to encode the latent representation $z_{bg}$, which captures common information.
 The decoder $p_{\theta}(x_{0}|x_{t},z_{bg})$ receives the noisy sample $x_{t}$ jointly with the common feature $z_{bg}$ to reconstruct the normal version of the input image.
-The saliency map is obtained by $\Delta_{SM}=\left|x_{0}-\hat{x_{0}}\right|$, where high values indicate the salient patterns.
+The saliency map is obtained by $\Delta_{SM}=\left|x_{0}-\hat{x_{0}}\right|$, where high intensity values indicate the anomaly patterns.
 Intuitively, the class of the image can be determined based on the magnitude of $\bar{\Delta_{SM}}$.
-Despite substantial progress in image classification models, they still face two significant challenges: i) dependence on extensive amounts of labelled data; and ii) limited interpretability.
-These challenges are particularly critical in sectors such as healthcare, where explaining the decision process is essential for clinicians to trust the model outcome, and unhealthy samples are scarce and hard to obtain Patrício et al.
-(2023).
-Contrastive Analysis (CA) represents a promising solution to this problem Zou et al.
-(2013); Abid et al.
-(2018); Weinberger et al.
-(2022) by learning the fundamental generative factors that distinguish a target (TG) dataset from a background (BG) dataset (referred to as salient patterns) and are shared across both datasets (referred to as common patterns).
-However, state-of-the-art CA methodologies have several problems: i) they often result in blurred images when utilizing Variational Autoencoders (VAEs) or are susceptible to mode collapse and unstable training when using Generative Adversarial Networks (GANs) Schlegl et al.
-(2019); Carton et al.
-(2024); ii) they depend on the availability of both healthy and unhealthy samples during training; iii) they struggle to preserve common patterns of the original image during the reconstruction process.
-To address these problems, this work redefines the CA task by employing a self-supervised contrastive encoder (referred to as common features encoder) to learn a latent representation, using samples exclusively from the BG dataset during training, and approximating the salient variation factor by leveraging data augmentation techniques such as random cutout or Gaussian noise (e.g.
-presence of eyeglasses in CelebA or presence of tumor in brain MRIs).
-The problems of previous generative methods are addressed by leveraging Denoising Diffusion Probabilistic Models (DDPM) Ho et al.
-(2020) to learn the distribution of the BG dataset (control group) conditioned on the learned latent representation by the common features encoder.
-During inference, the diffusion model processes unseen samples from BG or TG distributions by substituting the salient patterns with common patterns learned during training.
-This enables the classification of the input image as background (normal) or target (anomalous) by assessing the magnitude of the reconstruction error.
-Extensive validation on facial imagery data demonstrates improved similarity between original and reconstructed images.
-Additionally, comparative analysis with state-of-the-art methods for salient pattern detection for brain MRI further supports the efficacy of our approach.
-Our major contributions are as follows: i) a methodology for learning shared common information between background and target distributions, thus allowing the generation of a healthy (normal) version of the input image by encoding only its common information; ii) a common features encoder capable of learning input representations that are both class-invariant and instance-aligned.
-iii) a comprehensive evaluation of our approach on a facial imagery dataset and three brain MRI datasets, encompassing both healthy (background) and tumor (target) images.
+Despite substantial progress in image classification models, they still face two significant challenges: i) dependence on extensive amounts of labeled data; and ii) limited interpretability.
+These challenges are particularly critical in sectors such as healthcare, where explaining the decision process is essential for clinicians to trust the model outcome, and unhealthy samples are scarce and hard to obtain (Patrício et al., 2023).
+Contrastive Analysis (CA) represents a promising solution to this problem (Zou et al., 2013; Abid et al., 2018; Weinberger et al., 2022) by learning the fundamental generative factors that distinguish a target (TG) dataset from a background (BG) dataset (referred to as anomalous patterns) and are shared across both datasets (referred to as common patterns).
+However, state-of-the-art CA methodologies have several limitations: i) they often result in blurred images when utilizing Variational Autoencoders (VAEs) or are susceptible to mode collapse and unstable training when using Generative Adversarial Networks (GANs) (Schlegl et al., 2019; Carton et al., 2024); ii) they depend on the availability of healthy (BG) and unhealthy (TG) samples during training; iii) they struggle to preserve common patterns of the original image during the reconstruction process.
+On the other hand, UAD has emerged as an established approach for modeling the distribution of healthy anatomy and identifying deviations from this distribution as potential anomalies, thus removing the dependence on annotated unhealthy (TG) samples.
+While Denoising Diffusion Probabilistic Models (DDPM) (Ho et al., 2020) are commonly adopted as generative models in UAD, their training process lacks explicit constraints to preserve common patterns during reconstruction, hindering the fidelity of normal (healthy) reconstructions.
+Throughout this paper, we use the terms CA and UAD interchangeably, when referring to the task of detecting anomalous patterns in images.
+In this work, we propose an unsupervised methodology for the task of anomaly detection in brain MRIs.
+Specifically, we build upon a contrastive-guided conditional diffusion model (Figure 1) to reconstruct a healthy version of the input image while preserving its common anatomical patterns.
+This is achieved by conditioning the denoising diffusion process on a target-invariant latent feature.
+To this end, we employ a self-supervised contrastive encoder (referred to as target-invariant encoder) trained exclusively on samples from the healthy (BG) dataset.
+To approximate potential anomalous variation factors, we apply data augmentation techniques such as random cutout and Gaussian noise, which simulate features like tumor presence in brain MRIs.
+During inference, the diffusion model processes unseen samples belonging to the healthy (BG) or unhealthy (TG) distributions by reconstructing the input image such that the anomalous patterns are replaced with learned common (healthy) patterns.
+By computing pixel-wise differences between the original image and its generated counterpart, high-intensity values can be used to i) classify the input image as background (normal) or target (anomalous), based on the magnitude of the reconstruction error; and ii) localize potential anomalies, as illustrated in Figure 1 (bottom-right).
+Proof-of-concept validation on facial imagery data demonstrates the utility of our method.
+Additionally, an extensive comparative analysis against state-of-the-art methods for UAD on four brain MRI benchmarks highlights the effectiveness of our approach, achieving top performance on the recent and challenging evaluation-only NOVA benchmark (Bercea et al., 2025).
+Our major contributions are as follows:
+  • We propose a methodology for learning common features between background (healthy) and target (unhealthy) distributions, enabling the generation of a healthy (normal) counterpart of the input image by encoding only common information;
+  • We introduce a target-invariant encoder that learns representations which are both class-invariant and instance-aligned, facilitating common features preservation during the generative process;
+  • We conduct a comprehensive evaluation of our approach on a facial imagery dataset and four brain MRI datasets, encompassing both healthy (background) and tumor (target) images, to validate the generalization of our approach;
+  • Our method achieves state-of-the-art performance in anomaly localization on the challenging NOVA benchmark.
 2 Related Work
-Early work in contrastive analysis (CA) relied on the use of Variational Autoencoders (VAEs) Baur et al.
-(2021); Behrendt et al.
-(2022); Louiset et al.
-(2023).
+Early work in contrastive analysis (CA) relied on the use of Variational Autoencoders (VAEs) (Baur et al., 2021; Behrendt et al., 2022; Louiset et al., 2023).
 A recent method proposed by Louiset et al.
@@
 However, the resulting reconstructions were often blurry, limiting their interpretability and utility.
-Alternative approaches Schlegl et al.
-(2019); Carton et al.
-(2024) leveraging Generative Adversarial Networks (GANs) have been explored, but they suffer from issues like mode collapse and unstable training.
-More recently, Diffusion Denoising Probabilistic Models Song et al.
-(2021); Preechakul et al.
-(2022) (DDPMs) have emerged as a promising alternative for high-quality image generation, addressing the drawbacks of both GANs and VAEs.
-Our work aligns closely with the setting of salient pattern detection.
-In the domain of brain MRI, Behrendt et al.
-(2024) introduced a patch-based diffusion model for discovering salient patterns in an unsupervised manner in brain MRIs.
+Alternative approaches (Schlegl et al., 2019; Carton et al., 2024) leveraging Generative Adversarial Networks (GANs) have been explored, but they suffer from issues like mode collapse and unstable training.
+More recently, Diffusion Denoising Probabilistic Models (Song et al., 2021; Preechakul et al., 2022) (DDPMs) have emerged as a promising alternative for high-quality image generation, addressing the drawbacks of both GANs and VAEs.
+Our work aligns closely with the setting of CA and UAD.
+In the context of brain MRI, several works have been proposed (Behrendt et al., 2024; Iqbal et al., 2023; Behrendt et al., 2025).
+For example,  Behrendt et al.
+(2024) introduced a patch-based diffusion model for UAD in brain MRIs.
 They divide the input image into predefined patches and apply noise to each patch individually in the forward pass.
@@
 One drawback of their work is the extensive duration required for inference.
-Inspired by their approach, Iqbal et al.
-(2023) introduced masking-based regularization to train diffusion models, exploring various masking techniques applied at different levels of image modeling and frequency modeling.
-Nevertheless, their method also leads to increased inference time.
-Despite the potential of these methods for reconstruction fidelity, existing approaches struggle to preserve detailed characteristics of brain structure.
-To tackle this, we propose conditioning the denoising process of diffusion models with a latent representation encoding information about the common patterns of the input image, thereby preserving common information in the reconstructed output.
+More recently, Behrendt et al.
+(2025) proposed to conditioning a DDPM to better capture intensity characteristics and domain shifts in brain MRIs during reconstruction, thereby improving segmentation performance.
+In contrast, our method incorporates contrastive learning during encoder training to explicitly enforce a target-invariant embedding space.
 3 Background and Notations
@@
 Reverse Decoder
-Since the reverse process $q(x_{t-1}|x_{t})$ is intractable, it is used a DNN to approximate the distribution $p_{\theta}(x_{t-1}|x_{t})$, where $\theta$ represents the weights and biases of the network.
+Since the reverse process $q(x_{t-1}|x_{t})$ is intractable, a DNN is used to approximate the distribution $p_{\theta}(x_{t-1}|x_{t})$, where $\theta$ represents the weights and biases of the network.
 The reverse process is then modeled using a Gaussian distribution of the form:
@@
 From $p_{\theta}(x_{1:T})$ in Equation 3, one can generate a sample $x_{t-1}$ from a sample $x_{t}$ via:
-$x_{t-1}=\sqrt{\alpha_{t}-1}\left(\frac{x_{t}-\sqrt{1-\alpha_{t}}\epsilon_{\theta}(x_{t},t)}{\sqrt{\alpha_{t}}}\right)+\sqrt{1-\alpha_{t-1}}\epsilon_{\theta}(x_{t},t)+\sigma_{t}\epsilon_{t},$ (4)
+[…]
 where $\epsilon_{t}\sim\mathcal{N}(0,I)$ is standard Gaussian noise.
@@
 (2020) reformulated the variational lower bound objective function and considered the objective of predicting the total noise component added to the original image to create the noisy image at a given step.
-The loss function is then given by the squared difference between the predicted noise $\epsilon_{\theta}(x_{t},t)$ and the actual noise $\epsilon_{t}$, for a given time step $t$, using a U-Net Ronneberger et al.
-(2015):
+The loss function is then given by the squared difference between the predicted noise $\epsilon_{\theta}(x_{t},t)$ and the actual noise $\epsilon_{t}$, for a given time step $t$, using a U-Net (Ronneberger et al., 2015):
 $\mathcal{L}_{diff}=\left\|\epsilon_{\theta}(x_{t},t)-\epsilon_{t}\right\|_{2}^{2}.$ (5)
 By setting $\sigma_{t}=0$ in Equation (4), the coefficient before the random noise $\epsilon_{t}$ becomes zero, resulting in the following deterministic process of generating a new sample $x_{t}$.
-This modification allows to speed up the sampling process without degrading the quality of the generated samples Bishop and Bishop (2024).
+This modification allows to speed up the sampling process without degrading the quality of the generated samples (Bishop & Bishop, 2024).
 3.2 Contrastive Learning
@@
 Contrasting positive pairs against negative ones is an idea that dates back to previous research (Hadsell et al., 2006; Oord et al., 2019; Tian et al., 2020) and has seen various applications in different tasks, such as face recognition (Schroff et al., 2015).
-Let $x\in\mathcal{X}$ be an anchor sample, $x^{+}$ a positive sample (wrt to the anchor), and $x^{-}$ a negative sample.
+Let $x\in\mathcal{X}$ be an anchor sample, $x^{+}$ a positive sample (wrt the anchor), and $x^{-}$ a negative sample.
 CL methods look for a parametric mapping function $f:\mathcal{X}\rightarrow\mathbb{S}^{d-1}$ that maps semantically similar samples close together in the representation space (i.e.
@@
 Once pre-trained, $f$ is fixed, and its representation is evaluated on a downstream task, such as classification, through linear evaluation on a test set.
-Depending on how positive and negative samples are defined, CL can be employed in self-supervised Chen et al.
-(2020) or supervised Khosla et al.
-(2020) settings.
+Depending on how positive and negative samples are defined, CL can be employed in self-supervised (Chen et al., 2020) or supervised (Khosla et al., 2020) settings.
 4 Method
-4.1 Common Features Encoder
-The first block of our proposed approach is represented by the common features encoder.
+Our method employs a target-invariant encoder to extract a latent representation capturing common features of the input image.
+A diffusion-based model is then conditioned on this latent representation to reconstruct a normal (healthy) counterpart of the input.
+The absolute difference between the reconstructed and original images reveals potential anomalies, where brighter areas indicate abnormalities.
+An overview of the proposed approach is presented in Figure 1.
+4.1 Target-Invariant Encoder
+The first block of our proposed approach is represented by the target-invariant encoder.
 This encoder has the goal of learning an input representation which is invariant to the target variable (e.g.
-presence of eyeglasses in facial images or tumor on brain MRIs) but retains the common information of the input sample.
-The rationale of this approach is that an invariant representation can allow us to correctly reconstruct a realistic normal version of the input image, as it only encodes its common information.
-To explain the formulation of our encoder, we introduce two definitions specifying the properties that ensure that the encoder preserves the common information of the image.
+presence of eyeglasses in facial images or tumor in brain MRIs) but retains the common information of the input sample.
+Figure 1 (top-right) presents a visualization of the embedding space after projecting both healthy and unhealthy samples from the trained encoder.
+The rationale of this approach is that an invariant representation can allow us to correctly reconstruct a realistic normal version of the input image, as it only encodes its common (healthy) information.
+To explain the formulation of our encoder, we introduce three definitions specifying the properties that ensure that the encoder preserves the common information of the image.
 Definition 4.1.
@@
 6 can be expressed in terms of cosine similarity [^1] As representations are normalized, i.e.
-$||f(x)||_{2}=1$, then $sim(f(x),f(x^{+}))=1-d(f(x),f(x^{+}))$ where $d$ is a L2-distance function.: $sim(f(x),f(x^{-}_{j}))-sim(f(x),f(x^{+}))\leq-\epsilon\quad\forall j$ which corresponds to the $\epsilon$-InfoNCE loss Barbano et al.
-(2023):
+$||f(x)||_{2}=1$, then $sim(f(x),f(x^{+}))=1-d(f(x),f(x^{+}))$ where $d$ is a L2-distance function.: $sim(f(x),f(x^{-}_{j}))-sim(f(x),f(x^{+}))\leq-\epsilon\quad\forall j$ which corresponds to the $\epsilon$-InfoNCE loss (Barbano et al., 2023):
 $\mathcal{L}_{\epsilon-InfoNCE}=-\log\left(\frac{\exp(s^{+})}{\exp(s^{+}-\epsilon)+\sum_{j}\exp(s_{j}^{-})}\right)$ (7)
-where $s^{+}$ and $s^{-}_{j}$ are shorthand notations for $sim(f(x),f(x^{+}))$ and $sim(f(x),f(x^{-}_{j})$ respectively.
-To obtain a sample $x^{+}$ of the same subject of $x$, if it is not available in the training data, it is possible to employ an augmentation scheme such as in SimCLR Chen et al.
-(2020), i.e.
+where $s^{+}$ and $s^{-}_{j}$ are shorthand notations for $sim(f(x),f(x^{+}))$ and $sim(f(x),f(x^{-}_{j}))$ respectively.
+To obtain a sample $x^{+}$ of the same subject of $x$, if it is not available in the training data, it is possible to employ an augmentation scheme such as in SimCLR (Chen et al., 2020), i.e.
 $x^{+}=t(x)$ where $t\sim\mathcal{T}$ is an augmentation operator sampled from a family of standard augmentation $\mathcal{T}$ (e.g.
@@
 This means that the alignment in the latent space will be performed between samples with a different target attribute, hence achieving invariance.
-In a SSL setting, to avoid the dependence on anomalous samples, we leverage data augmentation and data manipulation techniques Dufumier et al.
-(2023) for approximating the distribution of the target attribute.
-For example, the appearance of tumors in brain MRIs can be approximated by employing random cutout, or Gaussian noise Behrendt et al.
-(2024) (details in Appendix B.2).
+In a SSL setting, to avoid the dependence on anomalous samples, we leverage data augmentation and data manipulation techniques (Dufumier et al., 2023) for approximating the distribution of the target attribute.
+For example, the appearance of tumors in brain MRIs can be approximated by employing random cutout, or Gaussian noise (Behrendt et al., 2024) (more details in the supplementary material).
 Definition 4.3.
-(Common Features Encoder) An encoder $f$ preserves common patterns of the image if it is both instance-aligned and class-invariant.
-The above definitions provide the theoretical support of the contrastive learning approach used for training the common features encoder.
+(Target-Invariant Encoder) An encoder $f$ preserves common patterns of the image if it is both instance-aligned and class-invariant.
+The above definitions provide the theoretical support of the contrastive learning approach used for training the target-invariant encoder.
 Considering that the learning process of the encoder is based on a contrastive learning strategy that is instance-aligned and class-invariant, our encoder is expected to produce features which preserve common information of the input image, regardless of whether the image belongs to the background or target set.
 4.2 Conditional Diffusion-based Decoder
-Our conditional diffusion decoder $p_{\theta}(x_{t-1}|x_{t},z_{bg})$ takes as input the noisy sample $x_{t}$ and the common feature $z_{bg}\in\mathbb{R}^{1\times d}=f_{\theta}(x_{0})$, a non-spatial vector of dimension $d$ that encodes common patterns observed in the input sample, derived from the properties of the common features encoder $f_{\theta}$ (Section 4.1).
+Our conditional diffusion decoder $p_{\theta}(x_{t-1}|x_{t},z_{bg})$ takes as input the noisy sample $x_{t}$ and the common feature $z_{bg}\in\mathbb{R}^{1\times d}=f_{\theta}(x_{0})$, a non-spatial vector of dimension $d$ that encodes common patterns observed in the input sample, derived from the properties of the target-invariant encoder $f_{\theta}$ (Section 4.1).
 Our primary objective is to reconstruct the background version of the input image, preserving its common information.
 Hence, we deviate from the step-wise sampling process in Equation (3), which is typically employed to generate new images from noise $p_{\theta}(x_{T:1}),x_{T}\sim\mathcal{N}(0,I)$.
-Instead, we directly estimate the input image $\hat{x_{0}}\sim p_{\theta}(x_{0}|x_{t},z_{bg})$ given the noisy sample $x_{t}$ (more details in Algorithm 2 in Appendix A.2).
-This is achieved by revising Equation (2), following Song et al.
-(2021), enabling to prediction of the denoised observation, which is an estimation of $x_{0}$ given $x_{t}$:
+Instead, we directly estimate the input image $\hat{x_{0}}\sim p_{\theta}(x_{0}|x_{t},z_{bg})$ given the noisy sample $x_{t}$ (more details in Algorithm 2 in the supplementary material).
+This is achieved by revising Equation (2), following (Song et al., 2021), allowing prediction of the denoised observation, which is an estimation of $x_{0}$ given $x_{t}$:
 $g_{\theta}(x_{t},t,z_{bg}):=(x_{t}-\sqrt{1-\alpha_{t}}\cdot\epsilon_{\theta}(x_{t},t,z_{bg}))/\sqrt{\alpha_{t}}$ (8)
-Then, the model is trained (Algorithm 1 in Appendix A.1) by minimizing the following objective function Preechakul et al. (2022), which is a modified version of the MSE objective in Equation (5):
+The model is then trained (Algorithm 1 in the supplementary material) by minimizing the following objective function (Preechakul et al., 2022), which is a modified version of the MSE objective in Equation (5):
 $\mathcal{L}_{diff}=\sum_{t=1}^{T}\mathbb{E}_{x_{0},\epsilon_{t}}\left[\left\|\epsilon_{\theta}(x_{t},t,z_{bg})-\epsilon_{t}\right\|_{2}^{2}\right]$ (9)
-We utilize adaptive group normalization layers (AdaGN) to condition the U-Net, as introduced in prior works Dhariwal and Nichol (2021); Preechakul et al.
-(2022).
-AdaGN incorporates the timestep and conditional variable embedding into each residual block by applying channel-wise scaling and shifting on the normalized intermediate activations $h\in\mathbb{R}^{c\times h\times w}$: $\textrm{AdaGN}(h,t,z_{bg})=z_{i}(t_{s}\textrm{GroupNorm}(h)+t_{b})$, where $z_{i}\in\mathbb{R}^{c}=\textrm{Affine}(z_{bg})$ and $(t_{s},t_{b})\in\mathbb{R}^{2\times c}=MLP(\psi(t))$ is the output of a multilayer perceptron incorporating a sinusoidal encoding function $\psi$.
-Additional details are provided in Appendix A.
-5 Experiments
-5.1 Experimental Data
+We condition the U-Net using adaptive group normalization (AdaGN), which integrates timestep and conditioning embeddings into each residual block via channel-wise scaling and shifting of normalized activations.
+Further details are provided in the supplementary material.
+5 Experimental Setup
+5.1 Datasets
 Facial Images
-We conduct initial experiments on CelebA Liu et al.
-(2015), containing 202,599 facial images with diverse attributes.
+We first conduct proof-of-concept experiments on CelebA (Liu et al., 2015) dataset, containing 202,599 facial images with diverse attributes.
 We create a subset focusing on subjects wearing eyeglasses and those without accessories, resulting in 15,353 images divided into two distinct classes: 1) Eyeglasses (EG): Images with the ’Eyeglasses’ attribute and no other accessory-related attributes, used solely for evaluation as target images; 2) No Eyeglasses (NEG): Images without the ’Eyeglasses’ attribute or any accessory-related attributes, used for training as background images.
-More information on dataset partitions and pre-processing is provided in Appendices B.1 and B.2.
-Medical Images
-To evaluate our common features encoder on medical data, we focus on tumor detection in brain MRIs.
-For a comprehensive comparison, we utilize the IXI IXI dataset as a background reference for training, as done in previous studies Behrendt et al.
-(2024); Iqbal et al.
-(2023).
-Evaluation is conducted on the Multimodal Brain Tumor Segmentation Challenge 2021 (BraTS21) Menze et al.
-(2014); Bakas et al.
-(2017); Baid et al.
-(2021) and the multiple sclerosis dataset (MSLUB) Lesjak et al.
-(2018), featuring tumor and Multiple Sclerosis (MS) samples, respectively.
+Our target task with this dataset is to remove glasses from input images while preserving subject-common information.
+Facial images are chosen due to their ease of visual interpretation, as identity preservation can be readily assessed during reconstruction.
+More details on dataset partitions and preprocessing are provided in the supplementary material.
+Brain MRIs
+We evaluate our approach in brain MRIs for the task of tumor detection and localization.
+For a comprehensive comparison, we utilize the IXI [^2] https://brain-development.org/ixi-dataset/ dataset as a background (healthy) reference for training, as in previous studies (Behrendt et al., 2024, 2025).
+Evaluation is conducted in the Multimodal Brain Tumor Segmentation Challenge 2021 (BraTS21) (Menze et al., 2014; Bakas et al., 2017; Baid et al., 2021) and the multiple sclerosis dataset (MSLUB) (Lesjak et al., 2018), featuring tumor and Multiple Sclerosis (MS) samples, respectively.
+Furthermore, we evaluate our model on the recently evaluation-only NOVA (Bercea et al., 2025) dataset, comprising 906 brain MRI scans spanning 281 rare and diagnostically diverse pathologies.
 Notably, only T2-weighted images are used from all datasets.
-More information on preprocessing and data partitioning is provided in Appendix B.3.
-5.2 Experimental Setup
-Implementation Details
-We compare our approach with state-of-the-art diffusion-based methods on CelebA Liu et al.
-(2015), including Diff-AE Preechakul et al.
-(2022), SepVAE Louiset et al.
-(2023), and DDIM Dhariwal and Nichol (2021).
-To ensure fair comparison, we use the official implementation of these methods and train them on our dataset partitions.
-For MRI experiments, we solely utilize our common features encoder alongside DDPM Wyatt et al.
-(2022), $p$DDPM Behrendt et al.
-(2024), and $m$DDPM Iqbal et al.
-(2023).
-These methods adopt distinct noise and objective functions, indirectly comparable to DDIM in facial images.
-Notably, they use structured simplex noise instead of Gaussian noise, capturing MRI image frequency distribution better.
-Additionally, they employ an alternative objective function to directly minimize $l1$ error between input image $x_{0}$ and its reconstruction $\hat{x_{0}}$ (Equation (9)).
-During training, MRI volumes are processed slice-wise, uniformly sampled with replacement.
-At test time, full volume reconstruction is achieved by iterating over all slices [^2] Implementation code of our method will be publicly available..
-Evaluation Pipeline
-During inference, the salient pattern map is obtained by the absolute difference between the input image $x_{0}$ and the reconstructed image $\hat{x_{0}}$, where higher values indicate larger reconstruction errors.
-To enhance the quality of these maps, inspired by Behrendt et al.
-(2024), we apply a median filter with a kernel size of 5 for score smoothing, followed by a morphological erosion operation for 10 iterations.
-Subsequently, the map is binarized, and the threshold is fine-tuned on the validation set to maximize the average DICE score.
-Additionally, we calculate the average Area Under the Precision-Recall Curve (AUPRC) on the test set.
-5.3 Results
-Table 1: Reconstruction quality of methods trained on CelebA (training set) and tested on CelebA (test set). NEG - No Eyeglasses, EG - Eyeglasses. $\left|\Delta\right|=\left|NEG-EG\right|$.
+More details on data preprocessing partitioning are provided in the supplementary material.
+5.2 Implementation Details
+As target-invariant encoder, we use a ResNet-50 backbone with a fully connected layer and a 128-dimensional output.
+For brain MRI experiments, we adapt the Spark2D (Tian et al., 2023) framework to incorporate the $\epsilon$-InfoNCE loss.
+Notably, the encoder is pre-trained in the healthy training set.
+At the inference phase, we fine-tune it along with the denoising network.
+For the diffusion model, structured simplex noise (Wyatt et al., 2022) is applied.
+During training, timesteps $t\in[1,1000]$ are sampled uniformly, while inference uses a fixed value of $t=500$.
+The denoising network follows a U-Net architecture.
+Training runs for up to 1600 epochs on NVIDIA A40 GPUs with Adam optimizer (learning rate of 1e-4, and a batch size of 32).
+Code is publicly available at https://github.com/CristianoPatricio/unsupervised-contrastive-cond-diff.
+6 Results
+Table 1: Reconstruction quality of methods trained on CelebA (training set) and tested on CelebA (test set).
+NEG - No Eyeglasses, EG - Eyeglasses.
+$\left|\Delta\right|=\left|NEG-EG\right|$.
+RE - Random Erasing.
 Model SSIM $\uparrow$ MSE $\downarrow$ LPIPS $\downarrow$
 NEG EG $\left|\Delta\right|$ NEG EG $\left|\Delta\right|$ NEG EG $\left|\Delta\right|$
-Diff-AE (No $\textrm{x}_{T}$)  Preechakul et al. (2022) $0.6287$ $0.6079$ $0.0208$ 0.01040.0104{0.0104}0.0104 $0.0119$ $0.0015$ $0.0847$ 0.10410.1041{0.1041}0.1041 $0.0194$
-SepVAE Louiset et al. (2023) $0.5796$ $0.4773$ $0.1023$ $0.0153$ $0.0268$ $0.0115$ $0.2912$ $0.3322$ $0.0410$
-DDIM Dhariwal and Nichol (2021) 0.96900.9690{0.9690}0.9690 $0.6568$ 0.3122 $\underline{0.0004}$ $0.0095$ $0.0091$ 0.00770.0077{0.0077}0.0077 $0.1385$ 0.1308
+Diff-AE (Preechakul et al., 2022) $0.6287$ $0.6079$ $0.0208$ 0.01040.0104{0.0104}0.0104 $0.0119$ $0.0015$ $0.0847$ 0.10410.1041{0.1041}0.1041 $0.0194$
+SepVAE (Louiset et al., 2023) $0.5796$ $0.4773$ $0.1023$ $0.0153$ $0.0268$ $0.0115$ $0.2912$ $0.3322$ $0.0410$
+DDIM (Dhariwal & Nichol, 2021) 0.96900.9690{0.9690}0.9690 $0.6568$ 0.3122 $\underline{0.0004}$ $0.0095$ $0.0091$ 0.00770.0077{0.0077}0.0077 $0.1385$ 0.1308
 Ours (w/o RE) $\underline{0.9756}$ $0.6737$ $0.3019$ 0.0003 $0.0097$ $0.0094$ $\underline{0.0060}$ $0.1198$ $0.1138$
 Ours (w/ RE) 0.9763 $0.6701$ $\underline{0.3062}$ 0.0003 $0.0099$ 0.0096 0.0059 $0.1213$ $\underline{0.1154}$
-5.3.1 CelebA Dataset
-Table 1 presents the reconstruction quality results of our method alongside Diff-AE Preechakul et al.
-(2022), SepVAE Louiset et al.
-(2023), and DDIM Dhariwal and Nichol (2021).
-These models were trained on the CelebA dataset using only background samples (NEG) and evaluated on both background (NEG) and target (EG) images.
-Evaluation metrics include SSIM Wang et al.
-(2004) ($\uparrow$), LPIPS Zhang et al.
-(2018) ($\downarrow$), and MSE ($\downarrow$).
-Our method (w/ RE) achieves the highest SSIM (0.9763) and lowest MSE (0.0003) and LPIPS (0.0059) when reconstructing background examples (NEG).
-While our method (w/o RE) and DDIM produce similar results, they fail to capture common information and high-level details compared to our method (w/ RE), as visually demonstrated in Figure 2.
-Additionally, ablation studies investigate the reconstruction quality when different $t\in[1,T]$ with $T=1000$ are chosen for corrupting the input image to predict its denoised version (Figure 10 in Appendix D).
-Empirically, we find that $t=250$ strikes a balance between image quality and common information preservation.
+We begin by validating the effectiveness of our method on the CelebA dataset.
+Subsequently, we evaluated its performance in anomaly detection and localization tasks, comparing it with established state-of-the-art UAD methods.
+6.1 Proof-of-Concept with CelebA Dataset
+Table 1 presents the reconstruction quality results of our method against baseline models (Preechakul et al., 2022; Louiset et al., 2023; Dhariwal & Nichol, 2021).
+These models were trained in the CelebA dataset using only background samples (NEG) and evaluated in both background (NEG) and target (EG) images.
+Evaluation metrics include SSIM ($\uparrow$), LPIPS ($\downarrow$), and MSE ($\downarrow$).
+Our method (w/ RE) achieves the highest SSIM (0.9763) and the lowest MSE (0.0003) and LPIPS (0.0059) when reconstructing background examples (NEG).
+Although our method (w/o RE) and DDIM produce similar results, they fail to capture common information and high-level details compared to our method (w/ RE), as visually demonstrated in Figure 2.
+Furthermore, the supplementary material includes ablation studies analyzing reconstruction quality for different corruption timesteps $t\in[1,1000]$ and various values of $\epsilon$ in $\epsilon$-InfoNCE.
+Figure 2: Reconstruction results of CelebA images using different methods.
+In contrast to Diff-AE, SepVAE and DDIM, our method produces images with well-preserved common information from the input image.
 EG-NEG Classification Performance
 Table 1 provides metrics for distinguishing between the two classes (NEG vs.
-EG), with the magnitude $\left|\Delta\right|=\left|NEG-EG\right|$ indicating distinctiveness.
-Histogram plots are available in Figure 7 in Appendix C.
-Diff-AE Preechakul et al.
-(2022) and SepVAE Louiset et al.
-(2023) exhibit overlapping distributions across metrics, while our method demonstrates well-separated distributions, yielding the highest ROC AUC score (Figure 8 in Appendix C).
+EG), with the magnitude $\left|\Delta\right|=\left|NEG-EG\right|$ indicating distinctiveness (histogram plots are available in the supplementary material).
+Diff-AE and SepVAE exhibit overlapping distributions across metrics, while our method demonstrates well-separated distributions, yielding the highest ROC AUC score (see Figure C.3 in the supplementary material).
 Instance-Alignment
@@
 We conduct an ablation study on $\epsilon$ to observe its impact on identification accuracy using a $k$-NN classifier with $k=3$.
-Results depicted in Figure 3 (left) show that the highest accuracy is achieved when $\epsilon$ is chosen from the ranges $[0.5,2.5]$ with the RE transformation and $[0.15,1.0]$ without it.
+Results depicted in Figure 3 (right) show that the highest accuracy is achieved when $\epsilon$ is chosen from the ranges $[0.5,2.5]$ with the random erasing (RE) transformation and $[0.15,1.0]$ without it.
+Figure 3: Left: Classification accuracy (NEG vs. EG) for different values of $\epsilon$. Right: Instance classification (accuracy in %) for different values of $\epsilon$.
 Importance of Class-Invariance
@@
 Subsequently, a logistic regression classifier was trained on the extracted image latent features.
-Results in Figure 3 show lower target accuracy with RE, suggesting stronger common information preservation.
+Results in Figure 3 (left) show lower target accuracy with RE, suggesting stronger common information preservation.
 However, instance classification performs worse with RE, possibly due to built-in augmentations in contrastive learning.
-Implementation details and numerical results are available in Appendix D.2.
-OriginalDiff-AESepVAEDDIMOurs(w/o RE)Ours(w/ RE)OriginalDiff-AESepVAEDDIMOurs(w/o RE)Ours(w/ RE)Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
-Figure 2: Reconstruction results of CelebA images using different methods.
-In contrast to Diff-AE, SepVAE and DDIM, our method produces images with well-preserved common information from the input image.
-$\epsilon$$\epsilon$Refer to captionRefer to caption
-Figure 3: Left: Instance classification (accuracy in %) for different values of $\epsilon$. Right: Classification accuracy (NEG vs. EG) for different values of $\epsilon$.
-Table 2: Comparison of the evaluated models on brain MRI datasets.
-The best results are highlighted in bold, and the second-best results are underlined.
-* denotes that the results were obtained by reproducing the method.
-The coloured values between brackets indicate the improvement w.r.t.
-the correspondent method.
-The reported results for all metrics correspond to the mean $\pm$ standard deviation across the different folds.
+6.2 Application on Brain MRI Datasets
+Table 2 compares different UAD methods on three brain MRI datasets using evaluation metrics such as DICE score, AUPRC, and $\ell_{1}$ reconstruction error (calculated only on healthy IXI dataset).
+We reproduce results for DDPM (Wyatt et al., 2022), pDDPM (Behrendt et al., 2024), and cDDPM (Behrendt et al., 2025), while results for other methods are taken from Behrendt et al.
+(2024).
+Table 2: Comparison of evaluated models in brain MRI datasets.
+Best results are shown in bold, and second-best results are underlined.
+* indicate results obtained by reproducing the method.
+All metrics are reported as mean $\pm$ standard deviation across 5 folds.
+Dice score and AUPRC are reported in percentage (%).
 Model BraTS21 MSLUB IXI
-DICE (%)\;$\uparrow$ AUPRC (%)\;$\uparrow$ DICE (%)\;$\uparrow$ AUPRC (%)\;$\uparrow$ $l1(1e-3)\downarrow$
-AE  Baur et al. (2021) $32.87$±$\pm 1.25$ $31.07$±$\pm 1.75$ $7.10$±$\pm 0.68$ $5.58$±$\pm 0.26$ $30.55$±$\pm 0.27$
-VAE Baur et al. (2021) $31.11$±$\pm 1.50$ $28.80$±$\pm 1.92$ $6.89$±$\pm 0.09$ $5.00$±$\pm 0.40$ $31.28$±$\pm 0.71$
-SVAE Behrendt et al. (2022) $33.32$±$\pm 0.14$ $33.14$±$\pm 0.20$ $5.76$±$\pm 0.44$ $5.04$±$\pm 0.13$ $28.08$±$\pm 0.02$
-DAE Kascenas et al. (2022) $37.05$±$\pm 1.42$ 44.9944.99{44.99}44.99±$\pm 1.72$ $3.56$±$\pm 0.91$ $5.35$±$\pm 0.45$ $10.12$±$\pm 0.26$
-f-AnoGAN Schlegl et al. (2019) $24.16$±$\pm 2.94$ $22.05$±$\pm 3.05$ $4.18$±$\pm 1.18$ $4.01$±$\pm 0.90$ $45.30$±$\pm 2.98$
-DDPM* Wyatt et al. (2022) $39.25$±$\pm 1.01$ $47.79$±$\pm 1.28$ $5.43$±$\pm 1.71$ $7.25$±$\pm 0.79$ $14.10$±$\pm 1.64$
-pDDPM* Behrendt et al. (2024) $49.47$±$\pm 0.91$ $54.68$±$\pm 1.02$ $9.17$±$\pm 1.29$ $\underline{10.35}$±$\pm 0.80$ $11.31$±$\pm 0.91$
-mDDPM* Iqbal et al. (2023) $\underline{51.77}$±$\pm 0.18$ $\underline{57.49}$±$\pm 0.41$ $5.03$±$\pm 0.37$ $9.47$±$\pm 0.02$ 7.287.28{7.28}7.28±$\pm 0.00$
-Ours (Common Features Encoder $+$ DDPM) 39.25±$\pm 0.55$ (+0.00) 48.58±$\pm 1.06$ (+0.79) $5.39$±$\pm 1.13$ 7.29±$\pm 0.23$ (+0.04) 13.72±$\pm 0.56$ (+0.38)
-Ours (Common Features Encoder $+$ pDDPM) $49.36$±$\pm 0.53$ $54.60$±$\pm 0.60$ 10.48±$\pm 2.19$ (+1.31) 10.2610.26{10.26}10.26±$\pm 0.67$ 11.11±$\pm 0.01$ (+0.20)
-Ours (Common Features Encoder $+$ mDDPM) $51.69$±$\pm 0.17$ 57.58±$\pm 0.17$ (+0.09) 9.47±$\pm 0.18$ (+4.44) $9.44$±$\pm 0.48$ $7.47$±$\pm 0.00$
-Identity-conditional sampling
-To demonstrate the efficacy of our encoder in capturing background patterns from a query image $x_{0}\sim q(x_{0})$, we encode $x_{0}$ to derive its latent representation $z_{bg}=f_{\theta}(x_{0})$.
-Then, using the step-wise denoising process $p_{\theta}(x_{t-1}|x_{t},z_{bg})$, we generate new images sharing the common information of $x_{0}$.
-Examples are shown in Figure 9 in Appendix C.
-With our encoder (w/ RE), the generated images preserve most background features and facial pose.
-Notably, subjects’ eyes appear more natural and perceptible with RE, suggesting the encoder is class-invariant, focusing on common patterns between the query and generated images.
-5.3.2 Brain MRI Datasets
-Table 2 provides a comparison between different unsupervised salient pattern detection methods in brain MRI.
-Evaluation metrics include DICE score, AUPRC, and $l1$ reconstruction error on healthy data (IXI).
-We reproduced results for DDPM Wyatt et al.
-(2022), pDDPM Behrendt et al.
-(2024), and mDDPM Iqbal et al.
-(2023).
-Results for other methods are referenced from the work of Behrendt et al.
-Behrendt et al.
-(2024).
-Reconstruction Results
-As depicted in Table 2, incorporating our common features encoder into diffusion-based models results in slight improvements across all datasets.
-Particularly noteworthy is the enhancement observed over the base results for the IXI dataset when our encoder is utilized.
-This suggests that when reconstructing a background (healthy) sample, the generated image preserves most of the common patterns from the input image.
-Detailed reconstruction examples can be found in Figure 11 in Appendix D.
-Salient Pattern Detection
-Examples of saliency maps generated by various methods are illustrated in Figure 4.
-Notably, we observe reduced foreground noise in the maps generated by $m$DDPM compared to DDPM, particularly evident in regions representing healthy brain anatomy.
-Specifically, when comparing the maps of the image depicted in the bottom-right row, the combination of Ours+DDPM and Ours+$m$DDPM demonstrates a smaller reconstruction error.
-[…]
-Figure 4: Salient pattern maps generated by different methods. Lighter colors indicate larger anomalies.
+DICE $\uparrow$ AUPRC $\uparrow$ DICE $\uparrow$ AUPRC $\uparrow$ $\ell_{1}(1e-3)\downarrow$
+AE (Baur et al., 2021) $32.87$±$\pm 1.25$ $31.07$±$\pm 1.75$ $7.10$±$\pm 0.68$ $5.58$±$\pm 0.26$ $30.55$±$\pm 0.27$
+VAE (Baur et al., 2021) $31.11$±$\pm 1.50$ $28.80$±$\pm 1.92$ $6.89$±$\pm 0.09$ $5.00$±$\pm 0.40$ $31.28$±$\pm 0.71$
+SVAE (Behrendt et al., 2022) $33.32$±$\pm 0.14$ $33.14$±$\pm 0.20$ $5.76$±$\pm 0.44$ $5.04$±$\pm 0.13$ $28.08$±$\pm 0.02$
+DAE (Kascenas et al., 2022) $37.05$±$\pm 1.42$ 44.9944.99{44.99}44.99±$\pm 1.72$ $3.56$±$\pm 0.91$ $5.35$±$\pm 0.45$ $10.12$±$\pm 0.26$
+f-AnoGAN (Schlegl et al., 2019) $24.16$±$\pm 2.94$ $22.05$±$\pm 3.05$ $4.18$±$\pm 1.18$ $4.01$±$\pm 0.90$ $45.30$±$\pm 2.98$
+DDPM* (Wyatt et al., 2022) $39.25$±$\pm 1.01$ $47.79$±$\pm 1.28$ $5.43$±$\pm 1.71$ $7.25$±$\pm 0.79$ $14.10$±$\pm 1.64$
+pDDPM* (Behrendt et al., 2024) $49.47$±$\pm 0.91$ $54.68$±$\pm 1.02$ $9.17$±$\pm 1.29$ 10.35±$\pm 0.80$ $11.31$±$\pm 0.91$
+cDDPM* (Behrendt et al., 2025) $50.57$±$\pm 1.34$ 56.12±$\pm 1.78$ $\underline{9.88}$±$\pm 0.83$ $9.45$±$\pm 0.60$ 9.71±$\pm 0.47$
+Ours 50.97±$\pm 2.06$ $56.11$±$\pm 2.19$ 10.08±$\pm 1.52$ $\underline{9.65}$±$\pm 1.35$ $9.92$±$\pm 0.59$
+Reconstruction Quality
+To assess overall reconstruction quality, we use the held-out test set from the healthy IXI dataset and report the $\ell_{1}$-error, calculated as the average absolute difference between the reconstructed images and their corresponding inputs.
+As shown in Table 2, our approach achieves performance competitive with cDDPM and pDDPM, and significantly outperforms all other baselines.
+These results suggest that the model effectively preserves healthy anatomy from the input image during reconstruction.
+This can be attributed to the influence of the $\epsilon$-InfoNCE regularization (Eq.
+7) within the target-invariant encoder, which improves the representation space with regard to the anatomy.
+Tumor Segmentation
+To evaluate segmentation performance in the context of UAD, we use all unhealthy test sets and report two metrics: i) Dice score and ii) Area Under Precision–Recall Curve (AUPRC).
+As shown in Table 2, our method is competitive with the top-performing baseline, showing marginal improvements in terms of Dice score.
+We achieve the best results with $\epsilon=0.5$ in Equation 7.
+Additional results with other $\epsilon$ values are available in the supplementary material.
+Analogous to the experimental findings on the CelebA dataset, this model can be used to distinguish between healthy and unhealthy MRIs by analyzing the generated saliency maps, as shown in Figure 1 (bottom-right).
+For instance, if the difference between a reconstructed brain MRI and its original counterpart shows high-intensity values across large regions, this may indicate presence of tumor and can be flagged for further review by a radiologist.
+This strategy can be seen as a form of clinical triage, while also contributing to the model’s explainability by providing a rationale for its decisions.
+Additional examples of reconstructions are shown in the supplementary material.
+Anomaly Localization on NOVA Benchmark
+Anomaly localization involves identifying and detect abnormal regions within brain MRIs.
+This task holds significant clinical importance, as many medical errors arise from failing to detect a pathology altogether.
+The objective is to predict one or more bounding boxes per image that correspond to abnormal regions, using annotations provided by radiologists as ground truth.
+As shown in Table 3, our method outperforms all baseline approaches in almost all evaluation metrics.
+Specifically, we achieve these results by considering only T2-weighted images and selecting the top-$5$ predicted largest bounding-boxes.
+We empirically found that including all modalities available in the NOVA dataset degrades performance, which is expected, as the models were pre-trained only on T2-weighted images from IXI.
+Compared to pDDPM and cDDPM, our pre-training strategy, which is based on contrastive learning, appears to be more robust than simply using data augmentation.
+As illustrated in Figure 4, our model is capable of detecting larger and more accurately localized anomalies, while baseline methods tend to produce smaller and more fragmented predictions.
+For example, pDDPM underestimates lesion size, likely due to its patching strategy, which may rely on suboptimal patch sizes.
+These results highlight the effectiveness of our methodology, particularly given that NOVA is considered a challenging and realistic evaluation-only benchmark dataset.
+Additional ablation studies on modality choice and the selection of top-$N$ largest bounding boxes are provided in the supplementary material.
+Table 3: Localization performance on NOVA benchmark.
+Reported metrics include standard object detection measures (mean Average Precision at multiple IoU thresholds), detection accuracy (ACC50), number of true positives (TP30), and number of false positives (FP30).
+Model mAP30 mAP50 mAP50-95 ACC50 TP30 FP30 $\downarrow$
+DDPM (Wyatt et al., 2022) 17.50 8.23 2.75 7.36 63/394 1517
+pDDPM (Behrendt et al., 2024) 0.30 0.30 0.06 0.25 1/394 1445
+cDDPM (Behrendt et al., 2025) 11.51 5.19 1.51 4.72 44/394 1423
+Ours 18.57 9.41 3.33 9.29 71/394 1492
+Figure 4: Examples of model predictions for anomaly localization on NOVA.
+Compared to baseline models, our method produces more precise and spatially focused bounding boxes that better align with ground-truth annotations (green).
+Each row corresponds to a distinct pathology.
 7 Conclusion and Future Work
-In this work, we introduce a self-supervised contrastive encoder, termed the common features encoder, designed to learn a latent representation capturing common patterns from input images.
-We then use diffusion-based models conditioned on this representation to generate realistic images exclusively encoding these common patterns.
-The encoder is trained to ensure class invariance and instance alignment, enabling the effective capture of common information during the reconstruction process.
-Our encoder demonstrates notable effectiveness in the CelebA dataset, and when integrated into established diffusion models for brain MRI, it improves the baseline results by preserving more background (common) patterns in the reconstructed images.
-Furthermore, investigating alternative mechanisms to integrate the latent condition into diffusion model architecture is a promising future direction.
-The potential implications of our approach on clinical workflows are discussed in Appendix E.
-Appendix A Conditional Diffusion Model Architecture
-The contrastive-guided conditional diffusion model is based on the DDIM Dhariwal and Nichol [2021] model, whereas the incorporation of the condition $z_{bg}$ into the U-Net model is based on the Diff-AE Preechakul et al.
-[2022] model.
-We selected the hyperparameters according to our computational resources.
-Figure 5 illustrates the operations inside the ResBlock to incorporate both the timestep $t$ and the condition $z_{bg}$, where $MLP$ denotes an embedding layer applied to both the timestep embedding and the condition $z_{bg}$ to match the number of output channels of the previous layer.
-Refer to caption $MLP(\psi(t))$ $MLP(z_{bg})$ $MLP(\psi(t))$ $MLP(z_{bg})$
-Figure 5: Diagram architecture of our approach. The bottom image illustrates a Residual Block along with the internal operations to include the timestep $t$ and the condition $z_{bg}$.
-A.1 Model Training
-The algorithm for model training is detailed in Algorithm 1.
-Algorithm 1 Conditional Diffusion Learning
-Input: Training data $\mathcal{D}=\{x_{n}\}$, Noise schedule $\{\beta_{1},...,\beta_{T}\}$
-Output: Network parameters w
-for $t\in\{1,...,T\}$ do
-   $\alpha_{t}\leftarrow\prod_{\tau=1}^{t}(1-\beta_{\tau})$
-end for
-repeat
-   $x\sim\mathcal{D}$
-   $t\sim\{1,...,T\}$
-   $z_{bg}\leftarrow f_{\theta}(x)$
-   $\epsilon\sim\mathcal{N}(0,I)$
-   ${x}_{t}\leftarrow\sqrt{\alpha_{t}}{x}+\sqrt{1-\alpha_{t}}\epsilon$
-   $\mathcal{L}(\textbf{w})\leftarrow\mathcal{L}_{diff}$
-   Take optimizer step
-until converge
-return w
-A.2 Model Inference
-The algorithm for predicting the normal (healthy) version of the input image is detailed in Algorithm 2.
-Algorithm 2 Denoising Process for Normal Image Inference
-Input: Test data $\mathcal{D}=\{x_{n}\}$, Noise schedule $\{\beta_{1},...,\beta_{T}\}$, Trained denoising network $\epsilon_{\theta}(x_{t},t,z_{bg})$
-Output: Denoised observation $\hat{x_{0}}$
-for $t\in\{1,...,T\}$ do
-   $\alpha_{t}\leftarrow\prod_{\tau=1}^{t}(1-\beta_{\tau})$
-end for
-$t\sim\{1,...,T\}$
-$\epsilon\sim\mathcal{N}(0,I)$
-${x}_{t}\leftarrow\sqrt{\alpha_{t}}{x}+\sqrt{1-\alpha_{t}}\epsilon$
-$z_{bg}\leftarrow f_{\theta}(x)$
-$\hat{x_{0}}\leftarrow g_{\theta}(x_{t},t,z_{bg})$
-return $\hat{x_{0}}$
-Appendix B Data Preprocessing
-B.1 CelebA Dataset
-The images per each split were selected based on the available official partitions and filtered based on the attribute annotations Liu et al.
-[2015].
-We filtered the CelebA dataset to obtain images containing the attribute Eyeglasses and images without any attribute related to wearing accessories.
-Specifically, we obtain the images of subjects with Eyeglasses by searching for the images matching the following attributes: Eyeglasses $=1$, Wearing_Hat $=-1$, Wearing_Earrings $=-1$, Wearing_Lipstick $=-1$, Wearing_Necklace $=-1$, Wearing_Necktie $=-1$.
-Detailed statistics of the filtered CelebA, including the train/validation/test partitions, are presented in Table 3.
-Table 3: CelebA Dataset Statistics.
-Split Train Validation Test
-Class NEG EG NEG EG NEG EG -
-# 6488 5931 800 764 712 658 -
-Total 12419 1564 1370 15353
-B.2 Custom Random Erasing
-We modify the RandomErasing class Zhong et al.
-[2020] available in the torchvision.transforms to include a parameter $b$ that controls the level of random noise (governed by a Gaussian distribution $\mathcal{N}(0,1-b),b\in[0,1]$) added to the rectangle region with respect to its spatial location and aspect ratio, where $b=1$ means that no noise is added.
-Figure 6 illustrates the augmented versions of the images (from CelebA and IXI datasets) after applying our custom random-erasing during the encoder training.
-Refer to caption Refer to caption
-(a) CelebA (b) IXI
-Figure 6: (a) - Left column: anchor image.
-Right column: augmented image after applying our custom random-erasing transformation.
-(b) - Left column: anchor image.
-Right column: augmented image after applying our custom random-erasing transformation.
-B.3 MRI Datasets
-  • IXI IXI dataset comprises 560 pairs of T1 and T2-weighted brain MRI scans.
-From the 560 pairs, 158 samples are used for testing and the remaining data is divided into 5 folds of 358 training samples and 44 validation samples for cross-validation.
-  • BraTS21 Menze et al.
-[2014], Bakas et al.
-[2017], Baid et al.
-[2021] dataset comprises 1251 brain MRI scans of four different weightings (T1, T1-CE, T2, FLAIR).
-The data is split into an unhealthy validation set of 100 samples and an unhealthy test set of 1151 samples.
-  • MSLUB Lesjak et al.
-[2018] dataset comprises brain MRI scans of 30 patients with Multiple Sclerosis (MS).
-For each patient, T1, T2, and FLAIR-weighted scans are available.
-The data is split into an unhealthy validation set of 10 samples and an unhealthy test set of 20 samples.
-Following Behrendt et al.
-[2024], we utilized T2-weighted images from all datasets across our experiments, as well as performed some pre-processing steps involving brain registration, skull stripping and downsampling.
-Please refer to Behrendt et al.
-[2024] for more detailed information on this pre-processing steps.
-Appendix C Additional Results
-Figure 7 shows the histograms plots of class-label distributions amongst MSE, SSIM and LPIPS on CelebA for the evaluated methods (Diff-AE, SepVAE and Ours).
-(a) Diff-AE Preechakul et al. [2022]
-(b) SepVAE Louiset et al. [2023]
-(c) Ours (w/ RE)
-Figure 7: Histogram of MSE, SSIM, and LPIPS for CelebA for NEG and EG samples.
-The ROC curves on CelebA for MSE, SSIM and LPIPS are illustrated in Figure 8.
-Refer to caption Refer to caption Refer to caption
-(a) MSE (b) SSIM (c) LPIPS
-Figure 8: ROC curve on CelebA for (a) MSE, (b) SSIM and (c) LPIPS.
-Figure 9 depicts examples of images generated by DDIM using the setp-wise denoining process $p_{\theta}(x_{t-1}|x_{t},z_{bg})$, conditioned on the common latent feature $z_{bg}$.
-QueryOurs (w/ RE)Ours (w/o RE)Close-upRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
-Figure 9: Identity-conditional generation using the latent representation $z_{bg}$ of the query image.
-Incorporating random-erasing (RE) during encoder training encourages class-invariance (Section 4.1), as demonstrated by the close-up on the eyes region, where w/o RE the eyes appear uniformly occluded.
-Appendix D Ablation Studies
-D.1 Sampling Facial Images at Different Timesteps
-Figure 10 shows the results for three facial images.
-We noticed that when choosing a $t>350$, the image gradually loses background (common) information.
-Conversely, $t<100$ fails to effectively replace eyeglasses, especially if they are sunglasses occluding the eyes (subject of first row).
-However, for normal eyeglasses without occluding the eyes (subject of last row), $t=100$ proves to be an optimal choice.
-[…]
-Figure 10: Predicted $\textrm{x}_{0}$ at different time steps. The predicted image gradually looses details and is altered as the level of noise increases.
-D.2 Instance-alignment and Class-invariant of Common Features Encoder
-To assess the influence of $\epsilon$ on both subject identification accuracy and EG/NEG classification accuracy, we conducted an ablation study on $\epsilon$ to verify the instance-alignment and class-invariance properties of our encoder (Section 4.1).
-The results are reported in Table 4.
-Specifically, for instance classification, we employed a $k$-NN classifier ($k=3$) and calculated the cosine similarity between the latent representation of each test sample and the remaining test samples.
-For EG/NEG classification, a logistic regression classifier was trained on the extracted latent features of images to distinguish between EG and NEG examples.
-Table 4: Results (Accuracy in %) of Instance Classification and EG/NEG Classification tasks on CelebA.
-$\epsilon$ w/o RE w/ RE
-EG/NEG Class. (%) Instance Class. (%) ($3-NN$) EG/NEG Class. (%) Instance Class. (%) ($3-NN$)
-0.0 80.219 24.14 76.715 22.55
-0.25 84.599 24.73 74.307 22.80
-0.5 82.993 26.32 74.088 22.72
-1.0 83.431 25.48 72.336 23.72
-1.5 82.555 22.97 72.701 23.05
-2.0 84.453 23.47 72.190 23.72
-2.5 80.876 24.06 70.949 22.72
-3.0 80.657 23.72 71.679 20.54
-3.5 80.803 24.48 68.686 21.29
-4.0 78.978 22.13 69.270 21.71
-4.5 78.467 23.22 68.832 19.87
-5.0 77.883 21.71 67.299 20.28
-6.0 78.686 23.81 64.161 19.95
-7.0 75.036 22.05 62.263 19.20
-8.0 73.358 20.54 62.774 20.12
-Diff-AE Preechakul et al. [2022] Target Accuracy: 87.664
-D.3 Reconstructed Brain MRI images
-In Figure 11, the reconstruction version of brain MRIs are depicted for the evaluated methods.
-Notably, the combination Ours+$p$DDPM achieved the lowest reconstruction $l1$ error.
-Ablation results on $\epsilon$ are present in Table 5.
-[…]
-Figure 11: Reconstruction results of MRI images using different methods.
-Table 5: Ablation results on $\epsilon$ for the combination of our encoder with the evaluated models on brain MRI datasets.
-The best results are highlighted in bold, and the second-best results are underlined.
-* denotes that the results were obtained by reproducing the method.
-The coloured values between brackets indicate the improvement/decrease w.r.t.
-the correspondent method.
-The reported results for all metrics correspond to the mean $\pm$ standard deviation across the different folds.
-Model BraTS21 MSLUB IXI
-DICE (%)\;$\uparrow$ AUPRC (%)\;$\uparrow$ DICE (%)\;$\uparrow$ AUPRC (%)\;$\uparrow$ $l1(1e-3)\downarrow$
-AE  Baur et al. [2021] $32.87$±$\pm 1.25$ $31.07$±$\pm 1.75$ $7.10$±$\pm 0.68$ $5.58$±$\pm 0.26$ $30.55$±$\pm 0.27$
-VAE Baur et al. [2021] $31.11$±$\pm 1.50$ $28.80$±$\pm 1.92$ $6.89$±$\pm 0.09$ $5.00$±$\pm 0.40$ $31.28$±$\pm 0.71$
-SVAE Behrendt et al. [2022] $33.32$±$\pm 0.14$ $33.14$±$\pm 0.20$ $5.76$±$\pm 0.44$ $5.04$±$\pm 0.13$ $28.08$±$\pm 0.02$
-DAE Kascenas et al. [2022] $37.05$±$\pm 1.42$ 44.9944.99{44.99}44.99±$\pm 1.72$ $3.56$±$\pm 0.91$ $5.35$±$\pm 0.45$ $10.12$±$\pm 0.26$
-f-AnoGAN Schlegl et al. [2019] $24.16$±$\pm 2.94$ $22.05$±$\pm 3.05$ $4.18$±$\pm 1.18$ $4.01$±$\pm 0.90$ $45.30$±$\pm 2.98$
-DDPM* Wyatt et al. [2022] $39.25$±$\pm 1.01$ $47.79$±$\pm 1.28$ $5.43$±$\pm 1.71$ $7.25$±$\pm 0.79$ $14.10$±$\pm 1.64$
-pDDPM* Behrendt et al. [2024] $49.47$±$\pm 0.91$ $54.68$±$\pm 1.02$ $9.17$±$\pm 1.29$ 10.35±$\pm 0.80$ $11.31$±$\pm 0.91$
-mDDPM* Iqbal et al. [2023] 51.77±$\pm 0.18$ $57.49$±$\pm 0.41$ $5.03$±$\pm 0.37$ $9.47$±$\pm 0.02$ $\underline{7.28}$±$\pm 0.00$
-Ours (Common Features Encoder $+$ DDPM)
-$\epsilon=0.5$ $38.68$±$\pm 0.85$ (-0.57) $48.49$±$\pm 1.47$ (+0.70) $6.47$±$\pm 1.51$ (+1.04) $7.69$±$\pm 1.30$ (+0.44) $13.78$±$\pm 0.52$ (+0.32)
-$\epsilon=1.0$ $39.01$±$\pm 0.60$ (-0.24) $48.01$±$\pm 1.30$ (+0.22) $5.44$±$\pm 2.06$ (+0.01) $7.97$±$\pm 0.99$ (+0.72) $13.85$±$\pm 0.60$ (+0.25)
-$\epsilon=2.0$ $39.25$±$\pm 0.55$ (+0.00) $48.58$±$\pm 1.06$ (+0.79) $5.39$±$\pm 1.13$ (-0.04) $7.29$±$\pm 0.23$ (+0.04) $13.72$±$\pm 0.56$ (+0.38)
-$\epsilon=3.0$ $39.21$±$\pm 0.36$ (-0.04) $48.08$±$\pm 1.57$ (+0.29) $5.35$±$\pm 1.25$ (-0.08) $7.38$±$\pm 0.81$ (+0.13) $13.79$±$\pm 0.42$ (+0.31)
-Ours (Common Features Encoder $+$ pDDPM)
-$\epsilon=0.5$ $48.93$±$\pm 1.42$ (-0.54) $54.06$±$\pm 1.54$ (-0.62) 10.55±$\pm 1.42$ (+1.38) $9.94$±$\pm 0.92$ (-0.41) $11.15$±$\pm 0.01$ (+0.16)
-$\epsilon=1.0$ $49.26$±$\pm 0.74$ (-0.21) $54.65$±$\pm 0.86$ (-0.03) $9.29$±$\pm 2.25$ (+0.12) $9.90$±$\pm 1.22$ (-0.45) $11.13$±$\pm 0.02$ (+0.18)
-$\epsilon=2.0$ $48.84$±$\pm 0.61$ (-0.63) $53.96$±$\pm 0.75$ (-0.72) $9.62$±$\pm 0.97$ (+0.45) $9.71$±$\pm 0.44$ (-0.64) $11.12$±$\pm 0.02$ (+0.19)
-$\epsilon=3.0$ $49.36$±$\pm 0.53$ (-0.11) $54.60$±$\pm 0.60$ (-0.08) $\underline{10.48}$±$\pm 2.19$ (+1.31) $\underline{10.26}$±$\pm 0.67$ (-0.09) $11.11$±$\pm 0.01$ (+0.20)
-Ours (Common Features Encoder $+$ mDDPM)
-$\epsilon=0.5$ $\underline{51.76}$±$\pm 0.85$ (-0.01) $\underline{57.56}$±$\pm 0.96$ (+0.07) $8.66$±$\pm 1.40$ (+3.63) $9.34$±$\pm 1.24$ (-0.13) $7.36$±$\pm 0.00$ (-0.08)
-$\epsilon=1.0$ $51.66$±$\pm 0.04$ (-0.11) $57.27$±$\pm 0.03$ (-0.22) $8.69$±$\pm 0.33$ (+3.66) $9.51$±$\pm 0.15$ (+0.04) 7.24±$\pm 0.79$ (+0.04)
-$\epsilon=2.0$ $51.50$±$\pm 0.58$ (-0.27) $57.15$±$\pm 0.57$ (-0.34) $8.75$±$\pm 0.96$ (+3.72) $9.22$±$\pm 0.42$ (-0.25) $7.39$±$\pm 0.00$ (-0.11)
-$\epsilon=3.0$ $51.69$±$\pm 0.17$ (-0.08) 57.58±$\pm 0.17$ (+0.09) $9.47$±$\pm 0.18$ (+4.44) $9.44$±$\pm 0.48$ (-0.03) $7.47$±$\pm 0.00$ (-0.19)
-Appendix E Potential Impact on Clinical Workflow
-The ability to reconstruct normal representations of input images, regardless of the presence of abnormalities, holds considerable significance in the disease diagnosis process.
-Early detection of diseases is aided by comparing the original image with its reconstructed healthy counterpart.
-Discrepancies between the images may signify the existence of anomalies or abnormalities, such as tumors.
-The development of an AI tool designed to assist healthcare practitioners in promptly identifying diseases or abnormalities by comparing two images enables timely intervention and treatment.
-Additionally, providing clinicians with reconstructed healthy images for comparison purposes can contribute to the reduction of diagnostic errors by providing clearer visualization of abnormalities and assisting in the process of differential diagnosis.
+In this work, we proposed an unsupervised framework for anomaly detection in brain MRIs that integrates contrastive learning with conditional diffusion models.
+Our method reconstructs a healthy version of input images while preserving common patterns, enabling both accurate classification and precise and interpretable localization of anomalies.
+By relying exclusively on healthy data, our approach mitigates the reliance on scarce or poorly annotated unhealthy samples.
+Moreover, the use of saliency maps derived from reconstruction errors provides intuitive visual evidence that enhances model interpretability, addressing key challenges in clinical adoption.
+Our results demonstrate the effectiveness of the proposed method, which achieves competitive performance compared to existing UAD methods and ranks first in anomaly localization on the challenging NOVA benchmark.
+Nevertheless, our work has limitations and open challenges that are important to acknowledge.
+The exclusive use of 2D T2-weighted images during training may limit generalization to other structural modalities.
+Future work could explore integrating multimodal imaging data and investigate the model’s robustness across diverse pathological conditions.
+The potential implications of our approach on clinical workflows are discussed in the supplementary material.